### v1

Abstract
Although Large Language Models (LLMs) show promising solutions to automated code generation, they often produce insecure code that threatens software security. Current approaches (e.g., SafeCoder) to improve secure code generation suffer from limited and imbalanced datasets, reducing their effectiveness and generalizability. In this work, we present Secure-Instruct, a novel framework that automatically synthesizes high‑quality vulnerable and secure code examples, generates fine‑tuning instructions, and instruction‑tunes LLMs to align task description and secure code generation abilities.
We evaluate Secure-Instruct on four representative LLMs using two benchmarks: our own CWEBench and the existing CWEval. CWEBench comprises 93 scenarios on 44 CWEs, all without overlap with Secure-Instruct’s synthetic instruction‑tuning dataset, while CWEval covers 31 CWEs with 119 manually verified security‑critical tasks. We find that Secure-Instruct improves not only the security but also the functional correctness of the generated code. On CWEBench, Secure-Instruct substantially improves secure code generation, giving a 14.3% average increase in secure ratio over the pretrained models and outperforms SafeCoder by 7.6%. On CWEval, Secure-Instruct achieves a 14% increase for CodeLlama-7B and 5.8% for Mistral-7B in Func-Sec@1 over pretrained models, and surpasses SafeCoder by 15.8% and 6.8% respectively.
1 Introduction
Large Language Models (LLMs), trained on billions of parameters across extensive corpora, have significantly advanced various software engineering tasks, such as code generation (Wei et al., 2024; Luo et al., 2023; Lin et al., 2024) and program repair (Jin et al., 2023; Joshi et al., 2023). Prominent models such as ChatGPT (OpenAI, 2024), Llama (Touvron et al., 2023), and GitHub Copilot (GitHub, 2024) have recently gained widespread popularity for their effectiveness in these domains.
Despite the success of LLMs in generating functionally correct code, researchers have identified quality concerns in LLM-generated code. One of the most critical issues is the presence of security vulnerabilities. Pearce et al. (Pearce et al., 2022) revealed that approximately 40% of code snippets generated by GitHub Copilot contain vulnerabilities. Recent studies (Liu et al., 2024b; Sandoval et al., 2023) also found similar security issues in code generated by other LLMs. These findings suggest that, despite advances, vulnerabilities continue to pose a significant challenge in LLM-based code generation.
Several efforts have been made to enhance the secure code generation capabilities of LLMs. Prompt-based approaches (Hajipour et al., 2023; Wang et al., 2023; He and Vechev, 2023) design and experiment with different security-aware prompts to guide LLMs in generating secure code. However, prompt-based approaches struggle to scale and generalize across a wide variety of CWEs, leading to inconsistent improvements and generalization issues. Alternatively, fine-tuning LLMs (He et al., 2024; Li et al., 2024a, b) has been explored to improve LLMs’ secure code generation capabilities. Among all, SafeCoder (He et al., 2024) is the state-of-the-art (SOTA), which leverages real vulnerability fixes for fine-tuning LLMs and outperforms all prompt-based approaches.
Yet, techniques like SafeCoder face two major limitations. First, exploiting real-world vulnerabilities and their fixes is limited by scale due to their rarity. For instance, SafeCoder managed to collect only 465 fixing examples covering 23 different CWEs with 6 programming languages from 145 million commits. This not only highlights the inefficiency of the approach but also its limited scalability, as even an extensive dataset contains examples for only a small subset of CWEs. Second, the methodology relies on a non-trivial manual inspection in the post-processing phase to balance and clean the dataset. Specifically, it down-samples over-represented CWE cases and manually removes examples that were incorrectly collected by the automated pipeline. Consequently, only 465 high-quality examples were retained from the 1,211 initially collected instances.
To address the abovementioned challenges, we propose Secure-Instruct, a novel framework designed to synthesize instruction tuning datasets using LLMs from descriptive documentation of CWEs. Specifically, we utilize the CWE descriptions and code examples provided by the MITRE MITRE (2025) database as a foundational resource to prompt LLMs to generate diverse, high-quality, and secure code examples that LLMs can easily understand for later fine-tuning. Secure-Instruct integrates static security detection tools (SonarQube and CodeQL) to automatically ensure the high quality of LLM-synthesized datasets. In addition, inspired by recent works in instruction tuning (Longpre et al., 2023), our proposed Secure-Instruct leverages instruction tuning to better align LLMs with secure code generation. Secure-Instruct (1) bridges the gap between CWE documentation and practical secure code generation, (2) improves the performance of secure code generation in LLMs, (3) requires no manual efforts in curating high-quality fine-tuning datasets, and (4) can be easily extensible to work on new CWEs as it does not rely on open-source repositories for curating high-quality fine-tuning dataset for new CWEs.
Secure-Instruct experiments two synthetic data generation schemes: (1) vulnerable code driven and (2) secure-code driven. For (1), given a CWE description document, Secure-Instruct prompts an LLM to generate vulnerable code first, then continuously fixes the vulnerable code to achieve secure code, resulting in code pairs of (vul, secure). For (2), Secure-Instruct prompts an LLM to generate secure code practices directly (secure). Then, Secure-Instruct leverages LLMs to generate an instruction for each code snippet. Last, Secure-Instruct employs two instruction tuning strategies: one is the standard one Wei et al. (2024), and the other one is the secure fine-tuning approach proposed by He et al. (He et al., 2024). This approach uses a fine-tuning technique designed to learn secure coding patterns and assign penalties to vulnerable code snippets.
We evaluate Secure-Instruct on four representative LLMs (CodeLlama-7B, CodeGen2-7B, Mistral-7B, and StarCoder-1B), which are commonly used in prior works on LLM secure code generation (He and Vechev, 2023; He et al., 2024) for fair comparison and are suitable under limited computing resources. Our evaluation includes two benchmarks, CWEBench (94 security scenarios covering 44 CWEs), and CWEval Peng et al. (2025) (119 manually verified security-critical tasks covering 31 CWEs). Our evaluation shows that the LLM fine-tuned by Secure-Instruct generates more secure code (an average improvement of 14.3%) compared to the pre-trained LLMs. On CWEBench, Secure-Instruct outperforms the state-of-the-art secure code generation approach, SafeCoder (He et al., 2024) by a large margin, +7.6% in CWEBench. On CWEval, Secure-Instruct achieves a better performance with +14.3% of Func-Sec@1 (both secure and functionally correct). Compared with the SOTA approach (SafeCoder), Secure-Instruct achieved +11.3% improvement in Func-Sec@1.
Furthermore, we performed ablation studies to understand the impact of fine-tuning datasets and approaches. Results show that the secure performance between Secure-Instructsecure and Secure-Instruct(vul, secure) is similar for the four models. Finally, we find that Secure-Instruct not only preserves but also improves LLMs’ capabilities in generating functionally correct code. The Pass@1 accuracy improved by +10.1% on CodeLlama-7B, +2.9% on Mistral-7B, and +1% on CodeGen2-7B based on the HumanEval benchmark (Chen et al., 2021). These results highlight the effectiveness of Secure-Instruct in advancing both secure and functional code generation.
The main contributions of this work include:
  • We propose Secure-Instruct, a fully automated framework that synthesizes high-quality datasets from CWE documentation to instruction tune LLMs for secure code generation. Secure-Instruct is generalizable and scalable to a diverse set of CWEs and is completely free of manual effort.
  • We extend existing secure benchmarks to a total of 93 security-sensitive prompts covering 44 CWEs to provide a comprehensive evaluation on secure code generation.
  • We conduct a comprehensive evaluation for Secure-Instruct on four LLMs, and we take a comparative analysis of secure performance with the SOTA approach. Our approach not only significantly improves the secure code generation but also improves the functionality correctness evaluated on CWEval, HumanEval, and MBPP benchmarks.
  • We share the synthesized fine-tuning dataset, the replication code, and fine-tuned models’ weights at https://zenodo.org/records/15019488
2 Background and Related Work
Security Concerns of LLM-Generated Code. Ensuring program security remains a challenge in both industry and academia. While LLMs have demonstrated significant potential in automating code generation, they are often trained on unsanitized datasets from open-source repositories, which can result in the generation of insecure code. Figure 1 shows an example of vulnerable code generated by CodeLlama-7B. This vulnerable code (lines 10 and 11) demonstrates a CWE-78 (‘OS Command Injection’). ‘username’ comes directly from ‘argv[1]’, which is not sanitized or validated. An attacker could inject arbitrary commands by passing a specially crafted username as input.
Several studies (Sandoval et al., 2023; Siddiq and Santos, 2023; Gong et al., 2024; Tihanyi et al., 2025; Fu et al., 2023; Mohsin et al., 2024) have investigated the security vulnerabilities present in LLM-generated code. Pearce et al. (Pearce et al., 2022) evaluated GitHub Copilot, which is built on the OpenAI Codex model, and reported that approximately 40% of the generated code contained security vulnerabilities. These vulnerabilities are systematically classified under the Common Weakness Enumeration (CWE) (CWE-MIRTE, 2022), a widely adopted category system for security vulnerabilities. Khoury et al. (Khoury et al., 2023) examined ChatGPT in security-critical coding scenarios, discovering that the model produced insecure code in 16 cases, with only 7 cases corrected after iterative interaction with the model. Similarly, Perry et al. (Perry et al., 2023) observed that developers relying on AI-assisted tools are more likely to introduce security vulnerabilities.
1int main(int argc, char *argv[]) {
2 // Get the username from the command line
3 char *username = argv[1];
4 // Get the user info from the system
5 char command[100] = {0};
6 // Execute the command using the username as input
7 // ** The code generated below **
8 sprintf(command, "id %s", username);
9}
Figure 1: An example of vulnerable code generation. Line 8 is generated by CodeLlama-7B.
Figure 2: An overview of Secure-Instruct.
Improving LLMs for Better Code Generation. Acquiring an effective and high-quality fine-tuning dataset from real-world sources is often expensive and labor-intensive. To address this challenge, researchers have explored methods for synthesizing datasets using large language models (LLMs). Recent advancements have introduced synthetic data generation techniques during fine-tuning to address data acquisition challenges and enhance LLM performance on specific tasks (Babbar and Schölkopf, 2019; Long et al., 2024; Gilardi et al., 2023; Chang et al., 2024; Liu et al., 2024a; Abay et al., 2019).
Self-Instruct Wang et al. (2022) introduced an automated approach for instruction tuning dataset generation using LLM. Subsequent works extend this strategy to a broader range of models and tasks Taori et al. (2023); Xu et al. (2024), including code generation Chaudhary (2023); Luo et al. (2023); Wei et al. (2024). Evol-Instruct (Xu et al., 2024; Luo et al., 2023) uses ChatGPT to generate diverse instruction-following examples using carefully crafted prompts. Wei et al. Wei et al. (2024) proposed OSS-Instruct, a method that uses ChatGPT to generate diverse coding problems and solutions from seed code snippets from open-source code, to create a high-quality fine-tuning dataset to improve code generation accuracy.
Improving the Security of LLM Generated Code.
To address security problems in LLM-generated code, Li et al. Li et al. (2024b) proposed an approach to fine-tune one LLM by using the vulnerability-fixing commits. Hajipour et al. (Hajipour et al., 2024) introduced a technique to synthesize secure code examples guided by CodeQL (CodeQL, 2024) oracle. Another work (Li et al., 2024a) explored the effect of parameter-efficient fine-tuning techniques in secure code generation. He et al. (He and Vechev, 2023) introduced SVEN, an advanced prompting method, i.e., prefix-tuning, that uses property-specific continuous vectors (prefixes) to guide LLMs to generate secure code and reduce the likelihood of generating vulnerable code. The training of SVEN optimizes these continuous vectors using a carefully curated, high-quality dataset.
The most recent work, SafeCoder (He et al., 2024), proposes a security-focused instruction tuning approach that learns from vulnerable–fix code pairs mined from 145 million commits. Through semi-automated filtering, the authors curate a dataset of 465 high-quality code examples. Their findings show that the LLM fine-tuned on this data achieves state-of-the-art performance in secure code generation.
Despite improved results, SafeCoder has several limitations. First, the quality of code from open-source repositories is inconsistent, often involving complex dependencies with external libraries. As a result, many code snippets are unsuitable for fine-tuning, requiring significant manual effort to filter and curate appropriate samples. Second, it is challenging to obtain a sufficiently large fine-tuning dataset for each category of security issues (e.g., CWE). For example, the work (He et al., 2024) identified 71 code snippets for CWE-079 but only one for CWE-732, resulting in an imbalance that hinders the model’s ability to achieve robust security performance. Besides, this approach is expensive. They only collected 456 code snippets covering 23 CWEs from 145 million commits, making it difficult to expand to other CWEs. To address these limitations, we propose Secure-Instruct to automatically construct a more extensive and diverse dataset to further enhance the security of LLM-generated code.
3 The Methodology of Secure-Instruct
We introduce Secure-Instruct, a framework that leverages LLMs (e.g., ChatGPT and Claude) to synthesize fine-tuning data for instruction-tuning code language models (CLMs) to enhance secure code generation. Figure 2 shows an overview of Secure-Instruct. It adopts two data-generation schemes, secure-code driven (Secure-Instructsecure) and vulnerable-code driven (Secure-Instruct(vul, secure)), to generate high-quality secure training data with no human intervention (Section 3.2). It then uses LLMs to generate natural-language instructions that describe the functionality of each code example in the dataset. Finally, Secure-Instruct applies two instruction tuning strategies to produce security-enhanced LLMs (Section 3.3): standard instruction tuning (Longpre et al., 2023) and security-specialized instruction tuning (He et al., 2024).
3.1 CWE Seeds Collection
First, we collect CWE documents from the MITRE. The MITRE is an official website that serves as an authoritative resource for understanding and analyzing Common Weakness Enumerations (CWEs). Each CWE entry on the website includes a comprehensive description, highlighting the nature of the vulnerability, its potential impact, and common scenarios in which it arises. Additionally, the entries feature illustrative code examples that contrast vulnerable and secure implementations, offering both context and practical insights.
Figure 3 presents an example of CWE-78 (OS Command Injection) from the MITRE website. This example includes a detailed overview of the weakness, an explanation of its root causes, and multiple code snippets demonstrating how this vulnerability can manifest in real-world software. These examples not only facilitate a deeper understanding of the security flaws but also serve as valuable references for LLMs to generate secure code examples. In particular, Secure-Instruct leverages the information from the MITRE website for each CWE, namely 1) the category of CWE, 2) the overall description, and 3) code snippet examples with their explanations. In total, we collected 44 CWEs from MITRE.
Figure 3: An example of CWE-78 on the MITRE website.
3.2 Synthetic Data Generation By Prompting LLMs
Given seed CWE documents, Secure-Instruct employs two data generation strategies to prompt LLM for generating fine-tuning datasets, namely vulnerable-code driven (Secure-Instruct(vul, secure)) and secure-code driven (Secure-Instructsecure).
Secure-Instruct(vul, secure). As Figure 2 illustrates, vulnerable-code driven data generation includes four steps: (1) prompting an LLM (e.g., GPT 4o) to generate vulnerable code given one CWE information; (2) verifying that the LLM-generated code indeed contains the specified CWE; (3) prompting LLMs to fix the vulnerabilities and produce secure code; and (4) confirming the secure code is vulnerability-free. Vulnerable-code driven generation ensures the success of generating both vulnerable and secure code for specific CWEs. We employ two powerful LLMs for code generation, i.e., ChatGPT (gpt-4o-2024-08-06) and Claude (claude-3-5-sonnet-20241022), to generate intentionally vulnerable code snippets. Below, we show the prompt templates for generating and fixing vulnerable code.
Prompt Template for Vulnerable Code Generation of Secure-Instruct(vul, secure). Task: Please generate an easily understandable vulnerable code snippet based on the following description of CWE. (Please consider the diversity of the vulnerable code examples for each generation.) CWE: {<CWE_OVERALL_DESCRIPTION>} A Vulnerable Code Example: This is {<LANGUAGE>} language example.
{<CODE>}
Explanation of the Example: {<EXPLANATION>} Can you generate a vulnerable code example for {<TARGET_LANGUAGE>} language?
Prompt Template for Fixing Vulnerable Code of Secure-Instruct(vul, secure) Task: Please fix the vulnerable code based on the following description of CWE : CWE: {<CWE_OVERALL_DESCRIPTION>} The Vulnerable Code:
{<GENERATED_VULNERABLE_CODE>}
Can you FIX the code for {<LANGUAGE>} ? Please make sure the code is secure to the CWE mentioned above and runnable.
For each pair of “CWE-ID” and “programming language”, Secure-Instruct(vul, secure) takes both the CWE description and a randomly selected code example from the seed corpus and prompts each LLM to generate 10 vulnerable code snippets. In total, Secure-Instruct(vul, secure) collected 1,560 code snippets spanning 78 CWE-language pairs. Then, Secure-Instruct(vul, secure) utilizes static security analyzers to verify the vulnerabilities present in the generated code. Finally, we instruct the LLMs to fix the vulnerable code and verify that the vulnerabilities are gone using static security analyzers.
Since static analysis tools can produce both false positives and false negatives, we use two static security analyzers, CodeQL (CodeQL, 2024) and SonarQube (SonarQube, 2025), to complement each other in verification. A code snippet is considered vulnerable only if both tools produce identical detection results. However, while CodeQL supports all 78 CWE-language pairs, SonarQube only supports a subset of 41 pairs (i.e., 6 CWEs for C, 9 for Java, 12 for Python, 9 for JavaScript, 0 for Ruby, and 5 for Go). For the cases not supported by SonarQube, Secure-Instruct solely relies on CodeQL.
After verification of the 1,560 snippets by static analyzers, 475 code snippets were confirmed to be truly vulnerable. These verified snippets then served as input seeds for the LLMs to generate five fixes for each snippet. Finally, static analyzers are used again to verify whether the generated fixes were valid. In the end, Secure-Instruct(vul, secure) collected 856 pairs of vulnerable code and their corresponding fixed code, which cover 25 CWEs in 6 programming languages, i.e., 39 CWE-language pairs.
Secure-Instructsecure. Secure-Instruct(vul, secure)’s double verification process results in many LLM-generated code snippets being filtered out, e.g., LLMs may fail to generate vulnerable code or fix vulnerable code in many cases, leading to a limited set of secure code snippets. To source an even larger set of secure code snippets, we propose an alternative approach, Secure-Instructsecure. Instead of generating vulnerable code and then repairing it, we directly instruct LLMs to produce secure code using CWE documents as prompts. For each CWE-language pair, we generate 100 secure code snippets that pass verification by both CodeQL and SonarQube, ensuring they do not contain vulnerabilities.
3.3 Instruction Tuning LLMs for Secure Code Generation
Secure-Instruct utilizes LLMs to generate a problem description as an “instruction” for each of the fine-tuning datasets. Then Secure-Instruct employs two instruction tuning methods, i.e., 1). Standard instruction tuning and 2). Secure instruction tuning.
Generating Instruction. We follow the work (He et al., 2024) to construct the instruction for each code snippet. The simplified prompt template is shown below:
Create a single, very short (maximum two sentences) non-detailed functionality description that could be used as a prompt to generate either of the code snippets below… {<The_Non-vulnerable_code_snippet>}
After that, we format the fine-tuning datasets by putting the code snippet to the <<response>> tag and the generated instructions to the <<instruction>> tag, following a prior work (Taori et al., 2023).
Standard Instruction Tuning. This process fine-tunes an LLM on instruction-response pairs to improve its ability to follow natural language instructions. It uses supervised fine-tuning (SFT), where the model learns by minimizing loss over the provided dataset. The loss of standard instruction tuning is defined as Equation (1), where $y_{<i}=(y_{1},y_{2},\dots,y_{i-1})$ denotes the tokens before $y_{i}$. This negative log-likelihood loss penalizes the model when it assigns low probability to the correct next token. It has been adopted by many prior studies (Wei et al., 2024; Xu et al., 2024; Luo et al., 2023).
$\mathcal{L}^{\text{std}}(x,y)=-\sum_{i=1}^{|y|}\log P\!\left(y_{i}\mid y_{<i},x\right)$ (1)
Secure-Specialized Instruction Tuning. SafeCoder (He et al., 2024) proposed a specialized fine-tuning technique for secure code generation by leveraging pairs of vulnerable code and fixed code mined from open-sourced software repositories. Since we collected 856 pairs of vulnerable and from Secure-Instruct(vul, secure), we applied this instruction tuning approach to our dataset. As shown in Equation (2), the secure loss focuses on encouraging the model to produce secure implementations, where $d^{\text{sec}}_{i}$ is a binary mask that identifies whether the $i$-th token in $y$ is security-critical using diff between secure and vulnerable code snippets. This masking ensures that the model receives a strong positive learning signal specifically on tokens that contribute to secure coding practices. Complementary to the secure loss, the vulnerability loss, shown in Equation (3), discourages the model from generating insecure tokens where $d^{\text{vul}}_{i}$ marking vulnerability-inducing tokens. By penalizing the likelihood of these insecure tokens, the model is pushed away from vulnerable code generation. Together, $\mathcal{L}^{\text{sec}}$ provides a positive signal for secure tokens, while $\mathcal{L}^{\text{vul}}$ provides a negative signal against vulnerable ones, which enables SafeCoder to promote secure code generation. Prior research (He and Vechev, 2023) highlights a critical concern: exclusively using secure code examples for fine-tuning can lead to a decline in code functionality. To address this limitation, we follow the work (He et al., 2024) via the integration of a secure code dataset and a functional code dataset during fine-tuning. By using standard instruction tuning for the functional dataset and secure instruction tuning for the secure dataset to improve both functional and secure performance for LLMs. Specifically, we incorporate the functional code dataset from the Evol-Instruct dataset (Luo et al., 2023), a high-quality dataset designed to enhance the functional robustness and generalization of code generation models.
$\mathcal{L}^{\text{sec}}(x,y,d)=-\sum_{i=1}^{|y|}d_{i}^{\text{sec}}\cdot\log P\!\left(y_{i}\mid y_{<i},x\right)$ (2)
$\mathcal{L}^{\text{vul}}(x,y,d)=-\sum_{i=1}^{|y|}d_{i}^{\text{vul}}\cdot\log\big(1-P(y_{i}\mid y_{<i},x)\big)$ (3)
4 Evaluation Setup
In this section, we describe our experimental setup.
Subject Large Language Models. We evaluate Secure-Instruct to improve four representative small and medium-sized LLMs with 1 to 7 billion parameters: CodeGen2-7B (Nijkamp et al., 2023), CodeLlama-7B (Roziere et al., 2023), Mistral-7B (Jiang et al., 2023), and StarCoder-1B (Li et al., 2023). These models are specifically designed for code generation, offering a balance between computational efficiency and performance. These models are widely used in prior works on secure code generation, such as (He and Vechev, 2023; He et al., 2024; Wei et al., 2024) for a fair comparison.
- CodeGen2-7B is a language model specifically designed for program synthesis. Trained on the “The Stack” dataset (Kocetkov et al., 2022), a comprehensive collection of diverse code repositories, CodeGen2 excels at generating and completing code snippets across multiple programming languages.
- CodeLlama-7B is developed by Meta, builds upon the Llama 2 (Touvron et al., 2023) architecture and is further fine-tuned specifically for code generation.
- StarCoder-1B is trained on “The Stack 1.2” dataset, a collection of code data and natural language text from GitHub, including over 80 programming languages, and the text extracted from GitHub issues and commits.
- Mistral-7B is developed by Mistral AI. It adopted Group-query attention (GQA) and sliding window attention (SWA) to increase the speed and lower the costs. It outperforms other models of the same size on a variety of benchmarks.
Details on Fine-Tuning LLMs. We use the parameter-efficient fine-tuning technique, LoRA Hu et al. (2021), to fine-tune the 7B LLMs. LoRA fine-tuning reduces the GPU memory usage during fine-tuning and maintains a promising performance. For the 1B model, we apply full-weight fine-tuning. All fine-tuning experiments are conducted on five Nvidia RTX A6000 GPUs with 48GB of memory. For full-weight fine-tuning, we set the learning rate to 2e-5, while for LoRA fine-tuning, the learning rate is 3e-4. Each model is fine-tuned in 2 epochs. In LoRA fine-tuning, we configure the LoRA alpha to 16 to control the scaling of LoRA layers and set the dropout rate to 0.05.
Evaluation Benchmarks. We evaluate Secure-Instruct using two secure-code generation benchmarks: (1) CWEBench crafted and extended by us, and (2) CWEval (Peng et al., 2025), which enables validation of both functional correctness and security by test cases.
First, we craft CWEBench by extending prior benchmarks for secure code generation (Siddiq and Santos, 2022; Pearce et al., 2022; He et al., 2024). Such prior benchmarks are composed of security-specific scenarios and are sourced from three resources, i.e., MITRE documentation (CWE-MIRTE, 2022), CodeQL repository (CodeQL, 2024), and scenarios manually crafted by the authors. We first combined all three existing benchmarks (Siddiq and Santos, 2022; Pearce et al., 2022; He et al., 2024), removed duplicates, and excluded the scenarios derived from MITRE examples, as Secure-Instruct uses MITRE documents as seeds for LLMs to generate fine-tuning datasets. Then, we followed a similar process as prior work He et al. (2024) to design 36 new security-specific scenarios covering 27 CWEs by adapting examples from the CodeQL repository. In the end, our crafted CWEBench consists of 93 security-specific scenarios covering 44 CWEs in six languages.
Second, we include a recent benchmark CWEval (Peng et al., 2025), which uses test cases for automated verification of both functional correctness and security without relying on static security checkers. In addition, CWEval provides manual-crafted comprehensive specifications that describe complex user intentions and logic, which means it is a much closer step to real-world applications. In total, CWEval consists of 119 security scenarios covering 31 CWE types in five programming languages.
De-duplication between fine-tuning datasets and evaluation benchmarks. To prevent data leakage between the fine-tuning dataset and the two benchmark datasets, we performed de-duplication between training and evaluation following  Zhou et al. (2025). Specifically, we used MinHash with Locality-Sensitive Hashing (LSH) to detect potential duplicates. Code snippets were grouped based on similarity, and pairs with a Jaccard similarity above 0.7 were flagged as potential duplicates. After that we removed these flagged snippets from the fine-tuning dataset.
5 Research Question Results
In this section, we present our three RQs. For each RQ, we detail the motivation, approach, and results.
RQ1: To what extent can Secure-Instruct enhance the capabilities of LLMs to generate secure code?
Motivation. We assess LLMs’ capabilities to generate secure code after being fine-tuned by Secure-Instruct. Approach. For each of the four studied LLMs, we experimented with two settings of Secure-Instruct: (1) applying security-specialized instruction tuning on the dataset of pairs of vulnerable and secure code, i.e., from Secure-Instruct(vul, secure); and (2) applying standard instruction tuning on the secure code snippets from Secure-Instructsecure. Table 1 summarizes the statistics of the resulting fine-tuning datasets using Secure-Instruct.
For CWEBench, we define a Secure Ratio to measure LLM’s capability to generate secure code. Secure Ratio of each security scenario is the percentage of secure code among all the generated code. The secure ratio is defined in Equation (3) where $S$ denotes the set of all generated code snippets and $\text{Sec}(s)$ evaluates to true if snippet $s$ is identified as non-vulnerable according to a static analyzer. Thus, the secure ratio quantifies the proportion of generated code that is free from vulnerabilities. To ensure robust results, we generated 100 code snippets for each security-specific scenario. A piece of LLM-generated code is deemed secure code if CodeQL detects no vulnerabilities in the code.
$\text{Secure Ratio}=\frac{\left|\{s\in S\mid\text{Sec}(s)\}\right|}{|S|}$ (4)
CWEval comes with three metrics, namely Func@k, Sec@k and Func-Sec@k to evaluate the security and functionality of generated code. In this study, we set k as 1. Specifically, Func@1 measures the percentage of security scenarios whose top-1 generated code is functionally correct (i.e., passes all functional test cases). Sec@1 is defined as the percentage of security scenarios whose top-1 solution is secure (i.e., passes all security test cases). Func-Sec@1 concerns both functional correctness and security. Given $n$ generated candidates, the unbiased Pass@k is computed by Equation (5. It defines the unbiased Pass@$k$ metric, denoted as $\text{Metric}_{\theta}@k$, where $\theta\in\{\text{Sec},\text{Func},\text{Func-Sec}\}$. Given $n$ generated candidates, the metric computes the probability that at least one of the top-$k$ candidates satisfies the evaluation criterion $\theta$. Thus, $\text{Metric}_{\theta}@k$ calculate the percentage of generated outputs that pass test cases under the chosen evaluation criterion within the top-$k$ samples.
$\text{Metric}_{\theta}@k\;=\;1-\frac{\binom{n-c_{\theta}}{k}}{\binom{n}{k}}\qquad\theta\in\{\text{Sec},\text{Func},\text{Func-Sec}\}$ (5)
Attribute Secure-Instruct(vul, secure) Secure-Instructsecure
# of Code Snippets 856 15,600
# of Average LOCs 52.0 50.0
# of CWEs 25 44
# of CWE-language pairs 39 78
Table 1: The statistics of the fine-tuning datasets used in Secure-Instruct.
Model Lang. Pre-trained (%) Secure-Instruct (%)
(vul,secure) secure
CodeLlama- 7B C 841/1,900 (44.3) 1,106/1,900 (58.2) 1,206/1,900 (63.5)
Go 341/900 (37.9) 729/900 (81.0) 707/900 (78.6)
Java 511/1,400 (36.5) 632/1,400 (45.1) 759/1,400 (54.2)
JS 937/1,800 (52.1) 1,235/1,800 (68.6) 1,294/1,800 (71.9)
Python 1,172/2,200 (53.3) 1,410/2,200 (64.1) 1,542/2,200 (70.1)
Ruby 621/1,100 (56.5) 877/1,100 (79.7) 988/1,100 (89.8)
Total 4,423/9,300 (47.6) 5,989/9,300 (64.4) 6,496/9,300 (69.8)
CodeGen2- 7B C 875/1,900 (46.1) 963/1,900 (50.7) 956/1,900 (50.3)
Go 488/900 (54.2) 390/900 (43.3) 617/900 (68.6)
Java 446/1,400 (31.9) 489/1,400 (34.9) 513/1,400 (36.6)
JS 956/1,800 (53.1) 805/1,800 (44.7) 1,147/1,800 (63.7)
Python 1,190/2,200 (54.1) 1,269/2,200 (57.7) 1,388/2,200 (63.1)
Ruby 584/1,100 (53.1) 638/1,100 (58.0) 683/1,100 (62.1)
Total 4,539/9,300 (48.8) 4,554/9,300 (49.0) 5,304/9,300 (57.0)
Mistral- 7B C 1,019/1,900 (53.6) 1,142/1,900 (60.1) 1,142/1,900 (60.1)
Go 395/900 (43.9) 693/900 (77.0) 721/900 (80.1)
Java 482/1,400 (34.4) 454/1,400 (32.4) 704/1,400 (50.3)
JS 1,005/1,800 (55.8) 1,174/1,800 (65.2) 1,168/1,800 (64.9)
Python 1,198/2,200 (54.5) 1,390/2,200 (63.2) 1,431/2,200 (65.0)
Ruby 695/1,100 (63.2) 847/1,100 (77.0) 1,008/1,100 (91.6)
Total 4,794/9,300 (51.5) 5,700/9,300 (61.3) 6,174/9,300 (66.4)
StarCoder- 1B C 1,097/1,900 (57.7) 1,306/1,900 (68.7) 1,256/1,900 (66.1)
Go 513/900 (57.0) 689/900 (76.6) 738/900 (82.0)
Java 549/1,400 (39.2) 560/1,400 (40.0) 549/1,400 (39.2)
JS 1,100/1,800 (61.1) 1,083/1,800 (60.2) 1,259/1,800 (69.9)
Python 1,152/2,200 (52.4) 1,413/2,200 (64.2) 1,583/2,200 (72.0)
Ruby 739/1,100 (67.2) 849/1,100 (77.2) 890/1,100 (80.9)
Total 5,150/9,300 (55.4) 5,900/9,300 (63.4) 6,275/9,300 (67.5)
Table 2: The secure ratio for pre-trained and fine-tuned models using Secure-Instruct on the full evaluation scenarios, broken down by language.
Results. Table 2 presents the results on CWEBench using Secure-Instruct. Across all models, Secure-Instruct(vul, secure) and Secure-Instructsecure demonstrate significant improvements in secure code generation. Compared with the pre-trained models, Secure-Instruct achieved the highest secure ratio across all models and languages. Specifically, CodeLlama‑7B achieved the highest secure ratio among the four models when fine‑tuned with Secure-Instructsecure, improving by 22.2% (from 47.6% to 69.8%), compared to a 16.8% increase under Secure-Instruct(vul, secure). CodeGen2-7B is improved from 48.8% to 57.0% by Secure-Instructsecure and to 49.0% by Secure-Instruct(vul, secure). Mistral-7B achieved an improvement of 14.9% in secure ratio by Secure-Instructsecure. StarCoder-1B achieves a notable improvement, i.e., secure ratio rising from 55.4% to 67.5%. Across the four LLMs, Secure-Instruct is shown to improve them to generate more secure code, and Secure-Instructsecure shows consistently better results than Secure-Instruct(vul, secure).
Table 3 presents the results on CWEval. Secure-Instruct achieves the highest Func-Sec@1 across all the studied LLMs. Notably, CodeLlama-7B and Mistral-7B show significant improvements, with increases of 14.0% and 7.8%, respectively, compared to their pre-trained versions. Overall, the LLMs tuned by Secure-Instruct exhibit better secure code generation than their pretrained counterparts on both CWEBench and CWEval, across multiple security metrics. On CWEBench, Secure-Instructsecure surpasses Secure-Instruct(vul, secure) by an average of 5.7% in secure ratio. On CWEval, the two perform comparably in terms of Func@1, Sec@1, and Func-Sec@1.
Secure-Instruct(vul, secure) v.s. Secure-Instructsecure. As Secure-Instructsecure utilizes a much larger fine-tuning dataset (15.6K v.s. 856), it is not surprising that Secure-Instructsecure outperforms Secure-Instruct(vul, secure). To understand more about the comparison, we down‑sampled the Secure-Instructsecure dataset to 865 examples, while ensuring samples covering across all CWE-language, to match the size of the Secure-Instruct(vul, secure) dataset. Table 4 shows the comparison results across the four LLMs. Notably, the differences between the two are minimal, indicating essentially equivalent effectiveness. This shows that Secure-Instructsecure still generates highly effective synthetic data without being driven by vulnerable code first. It also indicates that there exist duplicates in the data by Secure-Instructsecure and opportunities for future improvement.
Summary of RQ1: Secure-Instructsecure significantly improves the studied LLMs in secure code generation, achieving an average increase of over 10%. When using tuning datasets of the same size, Secure-Instructsecure and Secure-Instruct(vul, secure) achieve comparable performance.
Model Metrics Pre-trained SafeCoder Secure-Instruct
(vul,secure) secure
CodeLlama- 7B Func@1 22.1% 14.9% 37.9% 34.9%
Sec@1 17.8% 17.4% 31.2% 33.7%
Func-Sec@1 8.9% 7.1% 22.9% 22.9%
CodeGen2- 7B Func@1 20.7% NA 20.0% 20.9%
Sec@1 17.9% NA 17.7% 16.4%
Func-Sec@1 7.4% NA 7.2% 7.5%
Mistral- 7B Func@1 30.1% 21.7% 34.4% 31.3%
Sec@1 20.4% 17.4% 25.5% 27.9%
Func-Sec@1 12.5% 11.5% 17.9% 18.3%
StarCoder- 1B Func@1 15.3% NA 13.8% 12.7%
Sec@1 16.2% NA 16.4% 15.5%
Func-Sec@1 6.1% NA 7.4% 7.4%
Table 3: Func@1, Sec@1 and Func-Sec@1 results on CWEval for the pre-trained LLMs, SafeCoder fine-tuned LLMs, and Secure-Instruct (ours) fine-tuned LLMs. NA means Model not released by SafeCoder.
Models Secure-Instruct(vul, secure) Secure-Instructsecure
CodeLlama-7B 5,989/9,300 (64.4) 5,758/9,300 (61.9)
CodeGen2-7B 4,554/9,300 (49.0) 4,562/9,300 (49.1)
Mistral-7B 5,700/9,300 (61.3) 5,770/9,300 (62.0)
StarCoder-1B 5,900/9,300 (63.4) 6,157/9,300 (66.2)
Table 4: The secure ratios of Secure-Instruct(vul, secure) and Secure-Instructsecure with the same size (865) of training dataset.
CWE CodeLlama-7B Mistral-7B
Pre-trained SafeCoder Secure-Instruct Pre-trained SafeCoder Secure-Instruct
Secure (Vul, Secure) Secure (Vul, Secure)
CWE-022 16.0 43.2 32.0 37.2 14.8 34.6 38.0 20.8
CWE-078 16.8 79.0 50.7 40.8 30.5 64.2 60.0 52.0
CWE-079 37.8 84.6 64.8 67.0 29.2 68.8 68.6 48.6
CWE-089 52.0 70.2 82.5 72.5 51.7 60.8 67.0 69.0
CWE-116 66.7 65.0 69.7 75.3 65.7 68.7 71.3 65.3
CWE-119 87.0 99.0 67.0 47.0 93.0 100.0 5.0 96.0
CWE-190 24.0 7.0 23.0 23.0 25.0 1.0 5.0 15.0
CWE-200 24.0 99.0 99.0 17.0 34.0 98.0 73.0 50.0
CWE-295 10.5 78.5 82.0 78.0 17.5 78.5 78.0 85.0
CWE-326 40.0 33.3 83.0 98.0 55.7 72.7 59.3 60.3
CWE-327 47.8 68.4 64.0 64.0 61.2 54.6 77.0 74.8
CWE-338 38.5 42.0 56.0 71.0 13.5 31.5 50.5 46.0
CWE-352 87.7 93.3 95.7 100.0 94.7 85.7 97.3 96.3
CWE-502 58.3 79.1 97.3 80.7 68.3 89.4 93.9 80.0
CWE-611 4.0 54.0 40.0 0.0 0.0 51.0 40.0 0.0
CWE-676 38.0 48.5 47.5 49.5 56.0 54.5 93.0 53.5
CWE-681 58.0 50.5 56.5 54.5 54.0 35.5 80.5 50.5
CWE-732 32.0 75.0 99.5 91.0 57.0 89.5 93.5 92.5
CWE-915 24.0 82.0 95.0 97.0 25.0 51.0 37.0 77.0
CWE-916 94.0 97.0 100.0 98.0 96.0 95.0 99.0 100.0
Total (above) 43.4 67.9 70.4 66.3 48.2 64.9 69.4 62.3
CWE-020 75.0 50.5 75.0 61.0 61.0 75.0 78.0 54.0
CWE-094 88.0 84.2 99.2 98.0 88.8 82.0 91.5 97.2
CWE-117 14.5 7.5 9.5 2.5 8.5 4.0 7.5 4.5
CWE-120 77.0 47.5 94.0 93.5 58.0 80.5 97.0 92.5
CWE-129 25.0 48.5 90.0 75.0 26.0 64.0 73.5 61.5
CWE-134 27.0 0.0 3.0 0.0 91.0 1.0 1.0 1.0
CWE-170 59.0 14.0 72.0 48.0 68.0 27.0 18.0 32.0
CWE-178 77.0 80.0 85.0 99.0 93.0 70.0 93.0 99.0
CWE-209 60.2 47.8 77.8 62.3 74.0 46.8 78.2 60.2
CWE-215 42.0 31.0 62.0 71.0 68.0 71.0 29.0 71.0
CWE-243 26.0 69.0 99.0 96.0 68.0 87.0 86.0 96.0
CWE-273 63.0 61.0 16.0 77.0 51.0 38.0 92.0 73.0
CWE-311 1.0 0.0 0.0 0.0 4.0 0.0 6.0 0.0
CWE-312 58.0 63.0 97.0 100.0 45.0 85.0 77.0 95.0
CWE-367 75.0 95.0 97.0 92.0 94.0 76.0 96.0 94.0
CWE-369 26.0 42.0 54.0 46.0 35.0 13.0 3.0 17.0
CWE-401 37.0 99.0 93.0 33.0 51.0 78.0 8.0 64.0
CWE-532 3.0 0.0 50.0 16.0 13.0 1.0 19.0 5.0
CWE-643 96.0 98.0 99.0 100.0 94.0 85.0 100.0 100.0
CWE-770 65.0 67.0 97.0 81.0 83.0 64.0 99.0 75.0
CWE-777 4.5 7.5 20.5 20.0 3.0 32.0 27.0 14.0
CWE-788 87.0 93.0 100.0 100.0 74.0 87.0 99.0 93.0
CWE-798 72.0 73.0 94.0 69.0 66.3 73.3 93.7 92.0
CWE-918 47.5 10.5 17.0 14.5 24.5 8.5 2.5 1.0
Total 47.6 60.6 69.8 64.4 51.5 60.4 66.4 61.3
Table 5: The results of secure ratio in CodeLlama-7B and Mistral-7B. The top section lists CWEs that are included in the fine-tuning dataset of SafeCoder, while the bottom lists CWEs that are not covered in the SafeCoder dataset.
RQ2: How is Secure-Instruct compared with the state-of-the-art (SOTA) approach in securing LLM-based code generation?
Motivation. SafeCoder (He et al., 2024) represents the state-of-the-art in secure LLM-based code generation. It employs an automated pipeline to mine vulnerable code and corresponding fixes from a large corpus of open-source commits (145 million), followed by human intervention for post-processing and validation. The final dataset consists of (instruction, vulnerable code, secure code) triples used to fine-tune LLMs. In contrast, our proposed approach, Secure-Instruct , is a fully automated technique with zero human intervention. Given these fundamental differences, it is important to compare the performance of Secure-Instruct against SafeCoder to assess the effectiveness and scalability of our method.
Approach. We further conducted a comparative analysis between Secure-Instructsecure and SafeCoder on CWEBench and CWEval. In particular, we used the two released models, i.e., CodeLlma-7B and Mistral-7B, which were already fine-tuned by SafeCoder. In CWEBench, there are 44 CWE types. However, the training dataset of SafeCoder does not include all of them. SafeCoder was fine-tuned on a dataset containing 23 CWEs, whereas Secure-Instruct was fine-tuned on a dataset covering 44 CWEs. There are 20 CWEs covered by both Secure-Instruct and SafeCoder. We used the same configuration as described in the SafeCoder paper, namely a temperature of 0.8 and 100 samples for each security-specific scenario.
Results. Table 3 shows that Secure-Instruct achieves higher Func-Sec@1 on CWEval than SafeCoder for both CodeLlama‑7B and Mistral‑7B. Notably, SafeCoder exhibits a performance drop on CWEval. For example, SafeCoder’s CodeLlama-7B achieves only 7.1% Func-Sec@1, which is lower than the pre-trained model. A similar decline is observed in SafeCoder’s Mistral-7B. These findings are consistent with the observations reported in the study Peng et al. (2025).
Table 5 breaks down the detailed results on CWEBench of CodeLlama-7B and Mistral-7B in their pre-trained, SafeCoder, and Secure-Instructsecure, on each CWE in our evaluation benchmark. The highest secure ratio for each CWE is highlighted in bold. In the scenarios whose CWEs are covered in both approaches (the top section of the table), our analysis shows that CodeLlama-7B fine-tuned with Secure-Instructsecure achieved the highest overall secure ratio (70.6%), outperforming SafeCoder by 2.3% in secure ratio, Mistral-7B fine-tuned with Secure-Instructsecure achieved 68.9% secure ratio, with a 3.7% higher than SafeCoder. In the generalization evaluation (i.e., with the full security-specific scenarios of our benchmark), Secure-Instructsecure demonstrates strong generalization capabilities across a broader range of CWEs for both CodeLlama-7B and Mistral-7B. Specifically, CodeLlama-7B with Secure-Instructsecure significantly outperforms SafeCoder, achieving a 9.2% improvement, while Mistral-7B achieves a 6.0% increase. At the CWE category level, CodeLlama‑7B fine‑tuned with Secure-Instructsecure achieves the highest or ties for highest secure ratio in 29 of 44 CWEs, outperforming both its pretrained baseline and the SafeCoder‑fine‑tuned variant. Similarly, Mistral‑7B with Secure-Instructsecure leads in 25 of 44 CWEs relative to its pretrained and SafeCoder‑fine‑tuned counterpart. These results demonstrate that Secure-Instructsecure significantly enhances secure code generation, outperforming the SOTA approach, SafeCoder.
To assess statistical significance, we conducted Fisher’s exact tests comparing the number of secure code snippets generated by the pre-trained model and the model fine-tuned with Secure-Instructsecure for each CWE. CWEs with statistical significance (p<<0.05) are highlighted in bold, such as CWE-089. Notably, more than half of the CWEs in Secure-Instructsecure exhibit statistical significance, further demonstrating the robustness of our approach in enhancing secure code generation.
Summary of RQ2: Safecoder encounters a degradation in CWEval benchmark. Secure-Instructsecure significantly outperforms SafeCoder with 9.2% and 6.0% improvements across CWEBench. Even for those CWEs that are part of SafeCoder’s fine-tuning dataset, Secure-Instructsecure achieves 2.3% and 3.7% higher secure ratios in CodeLlama-7B and Mistral-7B, respectively.
RQ3: Can the LLM fine-tuned by Secure-Instruct preserve the functional correctness in code generation?
Motivation. A critical concern in fine-tuning LLMs in secure code generation is understanding how fine-tuning impacts the functionality of the generated code. While fine-tuning with secure code examples aims to improve the security aspects of the output, it may compromise the model’s ability to produce functionally correct code.
Approach. The benchmark, CWEval, evaluates the functionality of generated code by test cases with the metric Func@1. Besides, we evaluate the fine-tuned models on HumanEval and MBPP, two widely used benchmarks for assessing code generation in Python. For each task of HumanEval and MBPP, we generate 40 samples using both pre-trained and fine-tuned models and assess their functionality using the Pass@k metric, including Pass@1 and Pass@10. Pass@k is a standard metric for evaluating code generation, which represents the probability that at least one of the $k$ generated candidates for a given problem is correct.
Results. Table 3 presents the Func@1 results for all models. All models, except for StarCoder-1B, have improvements after using Secure-Instruct. CodeLlama-7B has an over 10% increase in Func@1 compared with the pre-trained model. Table 6 presents the Pass@1 (Pass@10 is available in our replication package) on the HumanEval and MBPP benchmarks for pre-trained LLMs, fine-tuned LLMs by SafeCoder and Secure-Instruct. For CodeLlama‑7B, among the pretrained, SafeCoder‑fine‑tuned, Secure-Instruct(vul, secure)‑fine‑tuned, and Secure-Instructsecure‑fine‑tuned variants, Secure-Instructsecure-fine-tuned model achieves the highest Pass@1 on HumanEval at 35.6%. For Mistral‑7B, the Secure-Instruct(vul, secure)‑fine‑tuned model achieves the highest score on MBPP at 30.7%. Overall, most models show performance improvements on both datasets, except for StarCoder-1B. On CWEval, its Func@1 dropped from 15.3% to 13.8% with Secure-Instruct(vul, secure), and 12.7% with Secure-Instructsecure. On HumanEval, the Pass@1 score of StarCoder-1B dropped from 11.7% to 10.9% with Secure-Instructsecureand from 11.7% to 6.7% with Secure-Instruct(vul, secure). This decline may be due to the smaller size of StarCoder-1B compared to the other LLMs.
Summary of RQ3: Secure-Instruct enhances secure code generation without sacrificing functional correctness. Across diverse LLMs and benchmarks, Secure-Instruct consistently outperforms both pre-trained LLMs and those fine-tuned by SafeCoder in generating functionally correct code.
Model Benchmark Pre-trained SafeCoder Secure-Instruct
(vul,secure) secure
CodeLlama- 7B HumanEval 25.5% 32.0% 34.7% 35.6%
MBPP 28.3% 33.7% 36.0% 34.7%
CodeGen2- 7B HumanEval 15.3% NA 8.1% 16.3%
MBPP 20.4% NA 15.4% 22.8%
Mistral- 7B HumanEval 23.0% 31.3% 31.6% 25.9%
MBPP 24.7% 29.3% 30.7% 27.1%
StarCoder- 1B HumanEval 11.7% NA 6.7% 10.9%
MBPP 13.7% NA 12.2% 17.7%
Table 6: Pass@1 results on HumanEval and MBPP for the pre-trained LLMs, SafeCoder fine-tuned LLMs, and Secure-Instruct (ours) fine-tuned LLMs.
6 Discussions
The accuracy of CodeQL. In CWEBench, the ground-truth of security is based on CodeQL, which is a static tool and may produce false positives. To address this concern, we randomly sampled 291 code snippets labeled as secure by CodeQL from four models and manually verified their security. All sampled snippets were confirmed to be truly secure. Additionally, we used another benchmark, CWEval, which evaluates security through test cases, to further validate our results.
The similarity between Secure-Instruct’s fine-tuning datasets and our evaluation benchmark. To clarify data leaks (i.e., overlap between fine-tuning dataset and benchmark), we performed a similarity analysis. Specifically, we first extracted all the secure code examples from the CodeQL repository for the 44 CWEs of our benchmark. Since most scenarios of the 44 CWEs are created based on the CodeQL repository, such secure code can be considered as ground-truth code for conducting such a similarity analysis. We compare each LLM-generated secure code with the secure code examples from the CodeQL repository under the same CWE. We then calculated the cosine similarity of each pair using TF-IDF embeddings (Sparck Jones, 1972). As there exist multiple secure examples for one CWE in the CodeQL repository, we selected the highest score, as the similarity score for an LLM-synthesized secure code snippet.
Figure 6 depicts the average cosine similarity score for the data generated by Secure-Instructsecure is 0.17, while the average score for data generated by Secure-Instruct(vul, secure) is 0.23. Secure-Instructsecure’s values align with findings from previous work (Wei et al., 2024) on synthesizing fine-tuning data using LLMs, which reported similarity scores typically ranging between 0.1 and 0.2 in the previous synthesis techniques. In contrast, Secure-Instruct(vul, secure) exhibits a slightly elevated similarity score, exceeding this range. These results indicate that Secure-Instructsecure generates data with lower similarity compared to Secure-Instruct(vul, secure). Importantly, this result indicates that the improvements from Secure-Instruct are not simply due to the inclusion of data from the same distribution of the CodeQL repository. Instead, the generated data provides additional diversity for fine-tuning the dataset.
Figure 4: Cosine similarities between Secure-Instruct’s fine-tuning dataset and the secure code examples by CodeQL (i.e., our evaluation benchmark).
Diversity of Secure-Instruct’s synthetic datasets. We followed a similar procedure to compute the pair-wise similarity for Secure-Instruct(vul, secure) and Secure-Instructsecure using TF-IDF. The similarity of Secure-Instruct(vul, secure) is 0.17. The similarity of Secure-Instructsecure is 0.16. Results show that datasets from both approaches are basically diverse.
Low cost of Secure-Instruct in generating fine-tuned datasets. We show that Secure-Instruct is able to generate high-quality fine-tuning data with low monetary cost and no human intervention. For example, when using ChatGPT (GPT-4o-2024-08-06), Secure-Instruct costs \$114.09 USD for generating all secure code snippets. Similarly, for Claude, the cost is \$109.74 USD for generating all secure code snippets. Compared to manually curated datasets, this synthesis method is highly cost-efficient. The cost per CWE-language is \$2.87 USD. Compared to our Secure-Instruct, prior work SafeCoder (He et al., 2024) imposes challenges on a high-efficient data curation process, i.e., a sharp reduction from hundreds of millions of code commits on open source repositories on GitHub to a total of 25,000 repositories that can be compiled and scanned by CodeQL for automatically detecting vulnerability-fixing commits, finally to merely 1,200 pairs of vulnerable and secure code examples. In addition, SafeCoder relies on manual effort to identify 465 high-quality code samples from the 1,211 samples produced by the automated pipeline. In contrast, Secure-Instruct demonstrates significant advancements in efficiency by reducing the time and cost required to construct high-quality datasets while maintaining a strong focus on generating reliable and secure code samples.
CodeLlama- 7B CodeGen2- 7B Mistral- 7B StarCoder- 1B
Secure-Instruct(vul, secure) 64.4% 49.0% 61.3% 63.4%
Secure-Instructsecure 61.9% 49.1% 62.0% 66.2%
Table 7: The secure ratios of Secure-Instruct(vul, secure) and Secure-Instructsecure with the same size (865) of training dataset.
Secure-Instruct(vul, secure) v.s. Secure-Instructsecure. Secure-Instructsecure uses a much larger dataset than Secure-Instruct(vul, secure) (15.6K vs. 856), so its better performance is expected. To compare fairly, we downsampled Secure-Instructsecure to 865 examples, covering all CWE-language pairs. Table 7 shows the comparison results across the four LLMs. Notably, the differences between the two are minimal, indicating essentially equivalent effectiveness. This shows that Secure-Instructsecure still generates highly effective synthetic data without being driven by vulnerable code first.
7 Threats to Validity
Internal Threats. LLMs can produce different outputs for the same prompt. To prevent the huge variations, we set a large sample size to ensure the stabilization results. CWEBench relies on CodeQL to detect vulnerability. Such a static analyzer may produce false positives. Although we manually reviewed samples, misclassifications could still influence the conclusions. We performed de-duplication across both our fine-tuning dataset and evaluation benchmarks (CWEBench and CWEval). However, this process may fail to detect duplicates, leading to false negatives.
External Threats. We follow prior work and choose LLMs with smaller parameter sizes to make fine-tuning feasible under limited computational resources. Fine-tuning larger models, such as ChatGPT-4o, would incur significantly higher costs. As a result, we acknowledge that our findings may not fully generalize to larger LLMs, and we consider this a direction for future investigation. We include 44 distinct CWEs in this study, reflecting a broad range of vulnerabilities. However, numerous additional CWEs remain unaddressed in this study. Moreover, individual CWE can manifest in various real-world scenarios, further highlighting the complexity of comprehensive evaluation. To improve coverage, we expanded our security-sensitive prompts to incorporate additional CWEs beyond those included in earlier studies. Compared to prior work (Pearce et al., 2022; He and Vechev, 2023; He et al., 2024; Li et al., 2024a), our study evaluates a significantly greater number of CWEs, providing a more thorough assessment of generated code vulnerabilities. Future efforts should include broader CWEs. Additionally, we include six programming languages in our evaluation benchmark. While these languages are highly popular and widely adopted in software development, our findings may not generalize to all programming languages. Future work should expand the language set to strengthen the robustness of the results.
8 Conclusion
In this paper, we proposed Secure-Instruct, a novel framework to improve LLM for secure code generation by harnessing LLMs for generating instruction tuning datasets. Secure-Instruct overcomes the limitations of the state-of-the-art technique, particularly addressing dataset imbalance, inefficient data curation, and the significant manual effort required. Unlike existing approaches, Secure-Instruct can be easily adapted to support new CWEs. Additionally, we contribute CWEBench, the largest evaluation benchmark to date for secure code generation. Our evaluation shows that Secure-Instruct significantly improves the performance of four representative LLMs, achieving 14.3% more secure code on average on CWEBench and +5.3% Func-Sec@1 on average on CWEval. Compared to the state-of-the-art, Secure-Instruct demonstrates strong performance even with a smaller fine-tuning dataset, highlighting its scalability and efficiency. Finally, we show that LLMs fine-tuned with Secure-Instruct maintain comparable or improved functional correctness in generated code.
Statements and Declarations
Funding. This work was supported by the Fonds de recherche du Québec (Grant No. 2024-NOVA-346499) [^1] https://repertoire.frq.gouv.qc.ca/offres/detailOffres.do;jsessionid=C333738136C878333C20227CC5EB4E1B?methode=afficher, the Natural Sciences and Engineering Research Council of Canada (NSERC) Discovery Grant (RGPIN-2019-07007), and the NSERC CREATE Grant (555406-2021). We gratefully acknowledge their support.
Competing Interests. The authors have no financial or proprietary interests in any material discussed in this article.
Data, Materials, and/or Code Availability. The data and code used in this study are available at https://zenodo.org/records/15019488.
Authors’ Contributions. Junjie Li - Conceptualization, Methodology, Data collection, Analysis, Writing original draft. Fazle Rabbi - Material preparation, Data collection, Analysis. Bo Yang - Material preparation, Data collection, Analysis, Writing original draft. Jinqiu Yang - Conceptualization, Methodology, Writing review & editing, Supervision, Supervising experiment design and execution. Song Wang - Conceptualization, Methodology, Writing review & editing, Supervision, Supervising experiment design and execution.

### v2

@@
-
+Secure-Instruct: An Automated Pipeline for Synthesizing Instruction-Tuning Datasets Using LLMs for Secure Code Generation
 Abstract
 Although Large Language Models (LLMs) show promising solutions to automated code generation, they often produce insecure code that threatens software security.
-Current approaches (e.g., SafeCoder) to improve secure code generation suffer from limited and imbalanced datasets, reducing their effectiveness and generalizability.
-In this work, we present Secure-Instruct, a novel framework that automatically synthesizes high‑quality vulnerable and secure code examples, generates fine‑tuning instructions, and instruction‑tunes LLMs to align task description and secure code generation abilities.
-We evaluate Secure-Instruct on four representative LLMs using two benchmarks: our own CWEBench and the existing CWEval.
+Current approaches (e.g., SafeCoder) to improve secure code generation are limited by small, imbalanced instruction-tuning datasets.
+In this work, we present Secure-Instruct, a novel pipeline that automatically synthesizes high‑quality vulnerable and secure code examples and instruction-tunes LLMs to align task description and secure code generation abilities.
+We evaluate Secure-Instruct on four representative LLMs using two security-related benchmarks: our own CWEBench and the existing CWEval.
 CWEBench comprises 93 scenarios on 44 CWEs, all without overlap with Secure-Instruct’s synthetic instruction‑tuning dataset, while CWEval covers 31 CWEs with 119 manually verified security‑critical tasks.
-We find that Secure-Instruct improves not only the security but also the functional correctness of the generated code.
-On CWEBench, Secure-Instruct substantially improves secure code generation, giving a 14.3% average increase in secure ratio over the pretrained models and outperforms SafeCoder by 7.6%.
-On CWEval, Secure-Instruct achieves a 14% increase for CodeLlama-7B and 5.8% for Mistral-7B in Func-Sec@1 over pretrained models, and surpasses SafeCoder by 15.8% and 6.8% respectively.
+We find that Secure-Instruct improves both security and functional correctness in code generation.
+On CWEBench, Secure-Instruct substantially improves secure code generation, giving a 28.5% increase on average in secure ratio over the pre-trained models and outperforms SafeCoder by 12.6%.
+On CWEval, Secure-Instruct achieves an increase of 157.3% for CodeLlama-7B and 46.4% for Mistral-7B in Func-Sec@1 over pretrained models, and significantly outperforms SafeCoder.
 1 Introduction
-Large Language Models (LLMs), trained on billions of parameters across extensive corpora, have significantly advanced various software engineering tasks, such as code generation (Wei et al., 2024; Luo et al., 2023; Lin et al., 2024) and program repair (Jin et al., 2023; Joshi et al., 2023).
-Prominent models such as ChatGPT (OpenAI, 2024), Llama (Touvron et al., 2023), and GitHub Copilot (GitHub, 2024) have recently gained widespread popularity for their effectiveness in these domains.
-Despite the success of LLMs in generating functionally correct code, researchers have identified quality concerns in LLM-generated code.
-One of the most critical issues is the presence of security vulnerabilities.
+Large Language Models (LLMs), trained on billions of parameters across extensive corpora, have significantly advanced various software engineering tasks, such as code generation (wei2024magicoder; luo2023wizardcoder; soen101) and program repair (jin2023inferfix; joshi2023repair).
+Prominent models such as ChatGPT (openai2024chatgpt), Llama (touvron2023llama), and GitHub Copilot (github2024copilot) have recently gained widespread popularity for their effectiveness in these domains.
+Despite their success, LLMs often generate code with security vulnerabilities.
 Pearce et al.
-(Pearce et al., 2022) revealed that approximately 40% of code snippets generated by GitHub Copilot contain vulnerabilities.
-Recent studies (Liu et al., 2024b; Sandoval et al., 2023) also found similar security issues in code generated by other LLMs.
+(pearce2022asleep) revealed that approximately 40% of code snippets generated by GitHub Copilot contain vulnerabilities, and similar issues have been reported in other models (liu2024no; sandoval2023lost).
 These findings suggest that, despite advances, vulnerabilities continue to pose a significant challenge in LLM-based code generation.
-Several efforts have been made to enhance the secure code generation capabilities of LLMs.
-Prompt-based approaches (Hajipour et al., 2023; Wang et al., 2023; He and Vechev, 2023) design and experiment with different security-aware prompts to guide LLMs in generating secure code.
-However, prompt-based approaches struggle to scale and generalize across a wide variety of CWEs, leading to inconsistent improvements and generalization issues.
-Alternatively, fine-tuning LLMs (He et al., 2024; Li et al., 2024a, b) has been explored to improve LLMs’ secure code generation capabilities.
-Among all, SafeCoder (He et al., 2024) is the state-of-the-art (SOTA), which leverages real vulnerability fixes for fine-tuning LLMs and outperforms all prompt-based approaches.
+Several efforts (hajipour2023codelmsec; wang2023enhancing; he2023large; he2024instruction; li2024exploratory; li2024fine) have been made to enhance the secure code generation capabilities of LLMs.
+Among all, SafeCoder (he2024instruction) is the state-of-the-art (SOTA), which leverages real vulnerability fixes for instruction-tuning LLMs and outperforms other approaches.
 Yet, techniques like SafeCoder face two major limitations.
 First, exploiting real-world vulnerabilities and their fixes is limited by scale due to their rarity.
-For instance, SafeCoder managed to collect only 465 fixing examples covering 23 different CWEs with 6 programming languages from 145 million commits.
+For instance, SafeCoder managed to collect only 465 fixing examples covering 23 different Common Weakness Enumerations (CWEs) with 6 programming languages from 145 million commits.
 This not only highlights the inefficiency of the approach but also its limited scalability, as even an extensive dataset contains examples for only a small subset of CWEs.
@@
 Consequently, only 465 high-quality examples were retained from the 1,211 initially collected instances.
-To address the abovementioned challenges, we propose Secure-Instruct, a novel framework designed to synthesize instruction tuning datasets using LLMs from descriptive documentation of CWEs.
-Specifically, we utilize the CWE descriptions and code examples provided by the MITRE MITRE (2025) database as a foundational resource to prompt LLMs to generate diverse, high-quality, and secure code examples that LLMs can easily understand for later fine-tuning.
+To address the abovementioned challenges, we propose Secure-Instruct, a novel and automated pipeline designed to synthesize instruction tuning datasets from descriptive documentation of CWEs.
+Specifically, we utilize the CWE descriptions and code examples provided by the MITRE (mitre-cwe) database as a foundational resource to prompt LLMs to generate diverse, high-quality, and secure code examples that LLMs can easily understand for later instruction-tuning.
 Secure-Instruct integrates static security detection tools (SonarQube and CodeQL) to automatically ensure the high quality of LLM-synthesized datasets.
-In addition, inspired by recent works in instruction tuning (Longpre et al., 2023), our proposed Secure-Instruct leverages instruction tuning to better align LLMs with secure code generation.
-Secure-Instruct (1) bridges the gap between CWE documentation and practical secure code generation, (2) improves the performance of secure code generation in LLMs, (3) requires no manual efforts in curating high-quality fine-tuning datasets, and (4) can be easily extensible to work on new CWEs as it does not rely on open-source repositories for curating high-quality fine-tuning dataset for new CWEs.
-Secure-Instruct experiments two synthetic data generation schemes: (1) vulnerable code driven and (2) secure-code driven.
+In addition, inspired by recent works in instruction tuning (longpre2023flan), our proposed Secure-Instruct leverages instruction tuning to better align LLMs with secure code generation.
+Secure-Instruct (1) bridges the gap between CWE documentation and practical secure code generation, (2) improves the performance of secure code generation in LLMs, (3) requires no manual efforts in curating high-quality instruction-tuning datasets, and (4) can be easily extensible to work on new CWEs as it does not rely on open-source repositories for curating high-quality instruction-tuning dataset for new CWEs.
+Secure-Instruct experiments with two synthetic data generation schemes: (1) vulnerable-code driven and (2) secure-code driven.
 For (1), given a CWE description document, Secure-Instruct prompts an LLM to generate vulnerable code first, then continuously fixes the vulnerable code to achieve secure code, resulting in code pairs of (vul, secure).
@@
 Then, Secure-Instruct leverages LLMs to generate an instruction for each code snippet.
-Last, Secure-Instruct employs two instruction tuning strategies: one is the standard one Wei et al.
-(2024), and the other one is the secure fine-tuning approach proposed by He et al.
-(He et al., 2024).
-This approach uses a fine-tuning technique designed to learn secure coding patterns and assign penalties to vulnerable code snippets.
-We evaluate Secure-Instruct on four representative LLMs (CodeLlama-7B, CodeGen2-7B, Mistral-7B, and StarCoder-1B), which are commonly used in prior works on LLM secure code generation (He and Vechev, 2023; He et al., 2024) for fair comparison and are suitable under limited computing resources.
-Our evaluation includes two benchmarks, CWEBench (94 security scenarios covering 44 CWEs), and CWEval Peng et al.
-(2025) (119 manually verified security-critical tasks covering 31 CWEs).
-Our evaluation shows that the LLM fine-tuned by Secure-Instruct generates more secure code (an average improvement of 14.3%) compared to the pre-trained LLMs.
-On CWEBench, Secure-Instruct outperforms the state-of-the-art secure code generation approach, SafeCoder (He et al., 2024) by a large margin, +7.6% in CWEBench.
-On CWEval, Secure-Instruct achieves a better performance with +14.3% of Func-Sec@1 (both secure and functionally correct).
-Compared with the SOTA approach (SafeCoder), Secure-Instruct achieved +11.3% improvement in Func-Sec@1.
-Furthermore, we performed ablation studies to understand the impact of fine-tuning datasets and approaches.
-Results show that the secure performance between Secure-Instructsecure and Secure-Instruct(vul, secure) is similar for the four models.
-Finally, we find that Secure-Instruct not only preserves but also improves LLMs’ capabilities in generating functionally correct code.
-The Pass@1 accuracy improved by +10.1% on CodeLlama-7B, +2.9% on Mistral-7B, and +1% on CodeGen2-7B based on the HumanEval benchmark (Chen et al., 2021).
+Last, Secure-Instruct employs two instruction tuning strategies: one is the standard one (wei2024magicoder), and the other is the secure instruction-tuning approach proposed by He et al.
+(he2024instruction).
+This approach uses an instruction-tuning technique designed to learn secure coding patterns and assign penalties to vulnerable code snippets.
+We evaluate Secure-Instruct on four representative LLMs (CodeLlama-7B, CodeGen2-7B, Mistral-7B, and StarCoder-1B), which are commonly used in prior works on LLM secure code generation (he2023large; he2024instruction) for fair comparison and are suitable under limited computing resources.
+Our evaluation includes two benchmarks, CWEBench (94 security scenarios covering 44 CWEs), and CWEval (peng2025cweval) (119 manually verified security-critical tasks covering 31 CWEs).
+On CWEBench, Secure-Instruct outperforms the state-of-the-art secure code generation approach, SafeCoder (he2024instruction) by a large margin, with an average improvement of 12.6%.
+On CWEval, Secure-Instruct achieves better performance with an average 56.6% improvement of Func-Sec@1 (both secure and functionally correct) compared to pre-trained models, and it outperforms SafeCoder with an average improvement of 140.8%.
+Finally, it also improves functional correctness with +39.6% (CodeLlama-7B), +12.6% (Mistral-7B), and +6.5% (CodeGen2-7B) on HumanEval (chen2021codex).
 These results highlight the effectiveness of Secure-Instruct in advancing both secure and functional code generation.
-The main contributions of this work include:
-  • We propose Secure-Instruct, a fully automated framework that synthesizes high-quality datasets from CWE documentation to instruction tune LLMs for secure code generation.
+Contributions.
+1) We propose Secure-Instruct, a fully automated framework that synthesizes high-quality datasets from CWE documentation to instruction tune LLMs for secure code generation.
 Secure-Instruct is generalizable and scalable to a diverse set of CWEs and is completely free of manual effort.
-  • We extend existing secure benchmarks to a total of 93 security-sensitive prompts covering 44 CWEs to provide a comprehensive evaluation on secure code generation.
-  • We conduct a comprehensive evaluation for Secure-Instruct on four LLMs, and we take a comparative analysis of secure performance with the SOTA approach.
+2) We extend existing secure benchmarks to a total of 93 security-sensitive prompts covering 44 CWEs to provide a comprehensive evaluation on secure code generation.
+3) We conduct a comprehensive evaluation for Secure-Instruct on four LLMs, and we take a comparative analysis of secure performance with the SOTA approach.
 Our approach not only significantly improves the secure code generation but also improves the functionality correctness evaluated on CWEval, HumanEval, and MBPP benchmarks.
-  • We share the synthesized fine-tuning dataset, the replication code, and fine-tuned models’ weights at https://zenodo.org/records/15019488
+4) We share the synthesized instruction-tuning dataset, the replication code, and models’ weights at https://zenodo.org/records/16715355.
 2 Background and Related Work
 Security Concerns of LLM-Generated Code.
-Ensuring program security remains a challenge in both industry and academia.
-While LLMs have demonstrated significant potential in automating code generation, they are often trained on unsanitized datasets from open-source repositories, which can result in the generation of insecure code.
+While LLMs show promise in automating code generation, their training on unsanitized open-source data often leads to insecure code.
 Figure 1 shows an example of vulnerable code generated by CodeLlama-7B.
-This vulnerable code (lines 10 and 11) demonstrates a CWE-78 (‘OS Command Injection’).
+This vulnerable code (line 8) demonstrates a CWE-78 (‘OS Command Injection’).
 ‘username’ comes directly from ‘argv[1]’, which is not sanitized or validated.
 An attacker could inject arbitrary commands by passing a specially crafted username as input.
-Several studies (Sandoval et al., 2023; Siddiq and Santos, 2023; Gong et al., 2024; Tihanyi et al., 2025; Fu et al., 2023; Mohsin et al., 2024) have investigated the security vulnerabilities present in LLM-generated code.
+Several studies (sandoval2023lost; siddiq2023generate; gong2024well; tihanyi2025secure; fu2023security; mohsin2024can) have investigated the security vulnerabilities present in LLM-generated code.
 Pearce et al.
-(Pearce et al., 2022) evaluated GitHub Copilot, which is built on the OpenAI Codex model, and reported that approximately 40% of the generated code contained security vulnerabilities.
-These vulnerabilities are systematically classified under the Common Weakness Enumeration (CWE) (CWE-MIRTE, 2022), a widely adopted category system for security vulnerabilities.
+(pearce2022asleep) evaluated GitHub Copilot, which is built on the OpenAI Codex model, and reported that approximately 40% of the generated code contained security vulnerabilities.
+These vulnerabilities are systematically classified under the Common Weakness Enumeration (CWE) (CWE), a widely adopted category system for security vulnerabilities.
 Khoury et al.
-(Khoury et al., 2023) examined ChatGPT in security-critical coding scenarios, discovering that the model produced insecure code in 16 cases, with only 7 cases corrected after iterative interaction with the model.
+(khoury2023secure) examined ChatGPT in security-critical coding scenarios, discovering that the model produced insecure code in 16 cases, with only 7 cases corrected after iterative interaction with the model.
 Similarly, Perry et al.
-(Perry et al., 2023) observed that developers relying on AI-assisted tools are more likely to introduce security vulnerabilities.
+(perry2023users) observed that developers relying on AI-assisted tools are more likely to introduce security vulnerabilities.
 1int main(int argc, char *argv[]) {
@@
 Figure 2: An overview of Secure-Instruct.
-Improving LLMs for Better Code Generation.
-Acquiring an effective and high-quality fine-tuning dataset from real-world sources is often expensive and labor-intensive.
-To address this challenge, researchers have explored methods for synthesizing datasets using large language models (LLMs).
-Recent advancements have introduced synthetic data generation techniques during fine-tuning to address data acquisition challenges and enhance LLM performance on specific tasks (Babbar and Schölkopf, 2019; Long et al., 2024; Gilardi et al., 2023; Chang et al., 2024; Liu et al., 2024a; Abay et al., 2019).
-Self-Instruct Wang et al.
-(2022) introduced an automated approach for instruction tuning dataset generation using LLM.
-Subsequent works extend this strategy to a broader range of models and tasks Taori et al.
-(2023); Xu et al.
-(2024), including code generation Chaudhary (2023); Luo et al.
-(2023); Wei et al.
-(2024).
-Evol-Instruct (Xu et al., 2024; Luo et al., 2023) uses ChatGPT to generate diverse instruction-following examples using carefully crafted prompts.
-Wei et al.
-Wei et al.
-(2024) proposed OSS-Instruct, a method that uses ChatGPT to generate diverse coding problems and solutions from seed code snippets from open-source code, to create a high-quality fine-tuning dataset to improve code generation accuracy.
 Improving the Security of LLM Generated Code.
-To address security problems in LLM-generated code, Li et al.
-Li et al.
-(2024b) proposed an approach to fine-tune one LLM by using the vulnerability-fixing commits.
+To overcome these limitations, Li et al.
+(li2024fine) proposed an approach to fine-tune one LLM by using the vulnerability-fixing commits.
 Hajipour et al.
-(Hajipour et al., 2024) introduced a technique to synthesize secure code examples guided by CodeQL (CodeQL, 2024) oracle.
-Another work (Li et al., 2024a) explored the effect of parameter-efficient fine-tuning techniques in secure code generation.
+(hajipour2024hexacoder) introduced a technique to synthesize secure code examples guided by CodeQL (codeql) oracle.
+Another work (li2024exploratory) explored the effect of parameter-efficient fine-tuning techniques in secure code generation.
 He et al.
-(He and Vechev, 2023) introduced SVEN, an advanced prompting method, i.e., prefix-tuning, that uses property-specific continuous vectors (prefixes) to guide LLMs to generate secure code and reduce the likelihood of generating vulnerable code.
+(he2023large) introduced SVEN, an advanced prompting method, i.e., prefix-tuning, that uses property-specific continuous vectors (prefixes) to guide LLMs to generate secure code and reduce the likelihood of generating vulnerable code.
 The training of SVEN optimizes these continuous vectors using a carefully curated, high-quality dataset.
-The most recent work, SafeCoder (He et al., 2024), proposes a security-focused instruction tuning approach that learns from vulnerable–fix code pairs mined from 145 million commits.
-Through semi-automated filtering, the authors curate a dataset of 465 high-quality code examples.
-Their findings show that the LLM fine-tuned on this data achieves state-of-the-art performance in secure code generation.
+Most recently, SafeCoder (he2024instruction) applies instruction-tuning on a dataset of 465 vulnerable-fix pairs, achieving state-of-the-art results in secure code generation.
 Despite improved results, SafeCoder has several limitations.
@@
 Second, it is challenging to obtain a sufficiently large fine-tuning dataset for each category of security issues (e.g., CWE).
-For example, the work (He et al., 2024) identified 71 code snippets for CWE-079 but only one for CWE-732, resulting in an imbalance that hinders the model’s ability to achieve robust security performance.
+For example, the work (he2024instruction) identified 71 code snippets for CWE-079 but only one for CWE-732, resulting in an imbalance that hinders the model’s ability to achieve robust security performance.
 Besides, this approach is expensive.
@@
 3 The Methodology of Secure-Instruct
-We introduce Secure-Instruct, a framework that leverages LLMs (e.g., ChatGPT and Claude) to synthesize fine-tuning data for instruction-tuning code language models (CLMs) to enhance secure code generation.
+We introduce Secure-Instruct that leverages LLMs (e.g., ChatGPT and Claude) to synthesize data for instruction-tuning code language models to enhance secure code generation.
 Figure 2 shows an overview of Secure-Instruct.
-It adopts two data-generation schemes, secure-code driven (Secure-Instructsecure) and vulnerable-code driven (Secure-Instruct(vul, secure)), to generate high-quality secure training data with no human intervention (Section 3.2).
+It adopts two data-generation schemes, secure-code driven (Secure-Instructsecure) and vulnerable-code driven (Secure-Instruct(vul, secure)), to generate high-quality secure training data with no human intervention.
 It then uses LLMs to generate natural-language instructions that describe the functionality of each code example in the dataset.
-Finally, Secure-Instruct applies two instruction tuning strategies to produce security-enhanced LLMs (Section 3.3): standard instruction tuning (Longpre et al., 2023) and security-specialized instruction tuning (He et al., 2024).
+Finally, Secure-Instruct applies two instruction-tuning strategies to produce security-enhanced LLMs (Section 3.3): standard instruction tuning (longpre2023flan) and security-specialized instruction tuning (he2024instruction).
 3.1 CWE Seeds Collection
-First, we collect CWE documents from the MITRE.
-The MITRE is an official website that serves as an authoritative resource for understanding and analyzing Common Weakness Enumerations (CWEs).
+First, we collect Common Weakness Enumeration (CWE) documents from the MITRE.
+The MITRE is an official website that serves as an authoritative resource for understanding and analyzing CWEs.
 Each CWE entry on the website includes a comprehensive description, highlighting the nature of the vulnerability, its potential impact, and common scenarios in which it arises.
 Additionally, the entries feature illustrative code examples that contrast vulnerable and secure implementations, offering both context and practical insights.
-Figure 3 presents an example of CWE-78 (OS Command Injection) from the MITRE website.
-This example includes a detailed overview of the weakness, an explanation of its root causes, and multiple code snippets demonstrating how this vulnerability can manifest in real-world software.
-These examples not only facilitate a deeper understanding of the security flaws but also serve as valuable references for LLMs to generate secure code examples.
-In particular, Secure-Instruct leverages the information from the MITRE website for each CWE, namely 1) the category of CWE, 2) the overall description, and 3) code snippet examples with their explanations.
+As shown in Figure 3, CWE-78 (OS Command Injection) provides such insights.
+Secure-Instruct leverages the information from the MITRE website for each CWE, namely 1) the category of CWE, 2) the overall description, and 3) code snippet examples with their explanations.
 In total, we collected 44 CWEs from MITRE.
@@
 3.2 Synthetic Data Generation By Prompting LLMs
-Given seed CWE documents, Secure-Instruct employs two data generation strategies to prompt LLM for generating fine-tuning datasets, namely vulnerable-code driven (Secure-Instruct(vul, secure)) and secure-code driven (Secure-Instructsecure).
+Given seed CWE documents, Secure-Instruct employs two data generation strategies to prompt LLM for generating instruction-tuning datasets, namely vulnerable-code driven (Secure-Instruct(vul, secure)) and secure-code driven (Secure-Instructsecure).
 Secure-Instruct(vul, secure).
-As Figure 2 illustrates, vulnerable-code driven data generation includes four steps: (1) prompting an LLM (e.g., GPT 4o) to generate vulnerable code given one CWE information; (2) verifying that the LLM-generated code indeed contains the specified CWE; (3) prompting LLMs to fix the vulnerabilities and produce secure code; and (4) confirming the secure code is vulnerability-free.
+As Figure 2 illustrates, the vulnerable-code driven data generation includes four steps: (1) prompting an LLM (e.g., GPT 4o) to generate vulnerable code given one CWE information; (2) verifying that the LLM-generated code indeed contains the specified CWE; (3) prompting LLMs to fix the vulnerabilities and produce secure code; and (4) confirming the secure code is vulnerability-free.
 Vulnerable-code driven generation ensures the success of generating both vulnerable and secure code for specific CWEs.
 We employ two powerful LLMs for code generation, i.e., ChatGPT (gpt-4o-2024-08-06) and Claude (claude-3-5-sonnet-20241022), to generate intentionally vulnerable code snippets.
-Below, we show the prompt templates for generating and fixing vulnerable code.
-Prompt Template for Vulnerable Code Generation of Secure-Instruct(vul, secure).
-Task: Please generate an easily understandable vulnerable code snippet based on the following description of CWE.
-(Please consider the diversity of the vulnerable code examples for each generation.) CWE: {<CWE_OVERALL_DESCRIPTION>} A Vulnerable Code Example: This is {<LANGUAGE>} language example.
-{<CODE>}
-Explanation of the Example: {<EXPLANATION>} Can you generate a vulnerable code example for {<TARGET_LANGUAGE>} language?
-Prompt Template for Fixing Vulnerable Code of Secure-Instruct(vul, secure) Task: Please fix the vulnerable code based on the following description of CWE : CWE: {<CWE_OVERALL_DESCRIPTION>} The Vulnerable Code:
-{<GENERATED_VULNERABLE_CODE>}
-Can you FIX the code for {<LANGUAGE>} ? Please make sure the code is secure to the CWE mentioned above and runnable.
-For each pair of “CWE-ID” and “programming language”, Secure-Instruct(vul, secure) takes both the CWE description and a randomly selected code example from the seed corpus and prompts each LLM to generate 10 vulnerable code snippets.
+The templates for generating and fixing vulnerable code are included in our replication package.
+For each pair of “CWE” and “programming language”, Secure-Instruct(vul, secure) takes both the CWE description and a randomly selected code example from the seed corpus and prompts each LLM to generate 10 vulnerable code snippets.
 In total, Secure-Instruct(vul, secure) collected 1,560 code snippets spanning 78 CWE-language pairs.
@@
 Finally, we instruct the LLMs to fix the vulnerable code and verify that the vulnerabilities are gone using static security analyzers.
-Since static analysis tools can produce both false positives and false negatives, we use two static security analyzers, CodeQL (CodeQL, 2024) and SonarQube (SonarQube, 2025), to complement each other in verification.
+Since static analysis tools can produce both false positives and false negatives, we use two static security analyzers, CodeQL (codeql) and SonarQube (sonarqube), to complement each other in verification.
 A code snippet is considered vulnerable only if both tools produce identical detection results.
-However, while CodeQL supports all 78 CWE-language pairs, SonarQube only supports a subset of 41 pairs (i.e., 6 CWEs for C, 9 for Java, 12 for Python, 9 for JavaScript, 0 for Ruby, and 5 for Go).
-For the cases not supported by SonarQube, Secure-Instruct solely relies on CodeQL.
-After verification of the 1,560 snippets by static analyzers, 475 code snippets were confirmed to be truly vulnerable.
+After verification of the 1,560 snippets by static analyzers, 475 were confirmed to be truly vulnerable.
 These verified snippets then served as input seeds for the LLMs to generate five fixes for each snippet.
 Finally, static analyzers are used again to verify whether the generated fixes were valid.
-In the end, Secure-Instruct(vul, secure) collected 856 pairs of vulnerable code and their corresponding fixed code, which cover 25 CWEs in 6 programming languages, i.e., 39 CWE-language pairs.
+After generating the dataset through our pipeline and verifying it using static analyzers, we performed a de-duplication step to eliminate potential overlaps with evaluation benchmarks.
+This ensures that our evaluation remains fair and unbiased.
+The details of this de-duplication process are described in Section 4.
+In the end, Secure-Instruct(vul, secure) produced 856 verified (vulnerable, fixed) pairs across 25 CWEs and 6 languages, covering 39 CWE-language combinations.
 Secure-Instructsecure.
@@
 For each CWE-language pair, we generate 100 secure code snippets that pass verification by both CodeQL and SonarQube, ensuring they do not contain vulnerabilities.
+Similarly, we performed a de-duplication step to make sure that all generated code snippets are not duplicated with evaluation benchmarks.
 3.3 Instruction Tuning LLMs for Secure Code Generation
-Secure-Instruct utilizes LLMs to generate a problem description as an “instruction” for each of the fine-tuning datasets.
+Secure-Instruct utilizes LLMs to generate a problem description as an “instruction” for each of the instruction-tuning datasets.
 Then Secure-Instruct employs two instruction tuning methods, i.e., 1).
@@
 Secure instruction tuning.
-Generating Instruction. We follow the work (He et al., 2024) to construct the instruction for each code snippet. The simplified prompt template is shown below:
-Create a single, very short (maximum two sentences) non-detailed functionality description that could be used as a prompt to generate either of the code snippets below… {<The_Non-vulnerable_code_snippet>}
-After that, we format the fine-tuning datasets by putting the code snippet to the <<response>> tag and the generated instructions to the <<instruction>> tag, following a prior work (Taori et al., 2023).
+Generating Instruction.
+We follow the work (he2024instruction) to construct the instruction for each code snippet using LLMs.
+After that, we format the instruction-tuning datasets by putting the code snippet in the <<response>> tag and the generated instructions in the <<instruction>> tag, following a prior work (alpaca).
 Standard Instruction Tuning.
 This process fine-tunes an LLM on instruction-response pairs to improve its ability to follow natural language instructions.
-It uses supervised fine-tuning (SFT), where the model learns by minimizing loss over the provided dataset.
-The loss of standard instruction tuning is defined as Equation (1), where $y_{<i}=(y_{1},y_{2},\dots,y_{i-1})$ denotes the tokens before $y_{i}$.
-This negative log-likelihood loss penalizes the model when it assigns low probability to the correct next token.
-It has been adopted by many prior studies (Wei et al., 2024; Xu et al., 2024; Luo et al., 2023).
-$\mathcal{L}^{\text{std}}(x,y)=-\sum_{i=1}^{|y|}\log P\!\left(y_{i}\mid y_{<i},x\right)$ (1)
+It uses supervised fine-tuning (SFT), where the model learns by minimizing loss over the provided dataset, which has been adopted by many prior studies (wei2024magicoder; xu2024wizardlm; luo2023wizardcoder).
 Secure-Specialized Instruction Tuning.
-SafeCoder (He et al., 2024) proposed a specialized fine-tuning technique for secure code generation by leveraging pairs of vulnerable code and fixed code mined from open-sourced software repositories.
+SafeCoder (he2024instruction) proposed a specialized instruction-tuning technique for secure code generation by leveraging pairs of vulnerable code and fixed code mined from open-sourced software repositories.
 Since we collected 856 pairs of vulnerable and from Secure-Instruct(vul, secure), we applied this instruction tuning approach to our dataset.
-As shown in Equation (2), the secure loss focuses on encouraging the model to produce secure implementations, where $d^{\text{sec}}_{i}$ is a binary mask that identifies whether the $i$-th token in $y$ is security-critical using diff between secure and vulnerable code snippets.
-This masking ensures that the model receives a strong positive learning signal specifically on tokens that contribute to secure coding practices.
-Complementary to the secure loss, the vulnerability loss, shown in Equation (3), discourages the model from generating insecure tokens where $d^{\text{vul}}_{i}$ marking vulnerability-inducing tokens.
-By penalizing the likelihood of these insecure tokens, the model is pushed away from vulnerable code generation.
-Together, $\mathcal{L}^{\text{sec}}$ provides a positive signal for secure tokens, while $\mathcal{L}^{\text{vul}}$ provides a negative signal against vulnerable ones, which enables SafeCoder to promote secure code generation.
-Prior research (He and Vechev, 2023) highlights a critical concern: exclusively using secure code examples for fine-tuning can lead to a decline in code functionality.
-To address this limitation, we follow the work (He et al., 2024) via the integration of a secure code dataset and a functional code dataset during fine-tuning.
-By using standard instruction tuning for the functional dataset and secure instruction tuning for the secure dataset to improve both functional and secure performance for LLMs.
-Specifically, we incorporate the functional code dataset from the Evol-Instruct dataset (Luo et al., 2023), a high-quality dataset designed to enhance the functional robustness and generalization of code generation models.
-$\mathcal{L}^{\text{sec}}(x,y,d)=-\sum_{i=1}^{|y|}d_{i}^{\text{sec}}\cdot\log P\!\left(y_{i}\mid y_{<i},x\right)$ (2)
-$\mathcal{L}^{\text{vul}}(x,y,d)=-\sum_{i=1}^{|y|}d_{i}^{\text{vul}}\cdot\log\big(1-P(y_{i}\mid y_{<i},x)\big)$ (3)
-4 Evaluation Setup
-In this section, we describe our experimental setup.
+Prior research (he2023large) highlights a critical concern: exclusively using secure code examples for instruction-tuning can lead to a decline in code functionality.
+To address this limitation, we follow the work (he2024instruction) via the integration of a secure code dataset and a functional code dataset during instruction-tuning.
+It uses standard instruction tuning for the functional dataset and secure instruction tuning for the secure dataset to improve both functional and secure performance for LLMs.
+Specifically, we incorporate the functional code dataset from the Evol-Instruct dataset (luo2023wizardcoder), a high-quality dataset designed to enhance the functional robustness and generalization of code generation models.
+4 Evaluation
+4.1 Evaluation Setup
 Subject Large Language Models.
-We evaluate Secure-Instruct to improve four representative small and medium-sized LLMs with 1 to 7 billion parameters: CodeGen2-7B (Nijkamp et al., 2023), CodeLlama-7B (Roziere et al., 2023), Mistral-7B (Jiang et al., 2023), and StarCoder-1B (Li et al., 2023).
+We evaluate Secure-Instruct to improve four representative small and medium-sized LLMs with 1 to 7 billion parameters: CodeGen2-7B (nijkamp2023codegen2), CodeLlama-7B (roziere2023code), Mistral-7B (jiang2023mistral), and StarCoder-1B (li2023starcoder).
 These models are specifically designed for code generation, offering a balance between computational efficiency and performance.
-These models are widely used in prior works on secure code generation, such as (He and Vechev, 2023; He et al., 2024; Wei et al., 2024) for a fair comparison.
-- CodeGen2-7B is a language model specifically designed for program synthesis.
-Trained on the “The Stack” dataset (Kocetkov et al., 2022), a comprehensive collection of diverse code repositories, CodeGen2 excels at generating and completing code snippets across multiple programming languages.
-- CodeLlama-7B is developed by Meta, builds upon the Llama 2 (Touvron et al., 2023) architecture and is further fine-tuned specifically for code generation.
-- StarCoder-1B is trained on “The Stack 1.2” dataset, a collection of code data and natural language text from GitHub, including over 80 programming languages, and the text extracted from GitHub issues and commits.
-- Mistral-7B is developed by Mistral AI.
-It adopted Group-query attention (GQA) and sliding window attention (SWA) to increase the speed and lower the costs.
-It outperforms other models of the same size on a variety of benchmarks.
-Details on Fine-Tuning LLMs.
-We use the parameter-efficient fine-tuning technique, LoRA Hu et al.
-(2021), to fine-tune the 7B LLMs.
-LoRA fine-tuning reduces the GPU memory usage during fine-tuning and maintains a promising performance.
-For the 1B model, we apply full-weight fine-tuning.
-All fine-tuning experiments are conducted on five Nvidia RTX A6000 GPUs with 48GB of memory.
-For full-weight fine-tuning, we set the learning rate to 2e-5, while for LoRA fine-tuning, the learning rate is 3e-4.
-Each model is fine-tuned in 2 epochs.
-In LoRA fine-tuning, we configure the LoRA alpha to 16 to control the scaling of LoRA layers and set the dropout rate to 0.05.
-Evaluation Benchmarks.
-We evaluate Secure-Instruct using two secure-code generation benchmarks: (1) CWEBench crafted and extended by us, and (2) CWEval (Peng et al., 2025), which enables validation of both functional correctness and security by test cases.
-First, we craft CWEBench by extending prior benchmarks for secure code generation (Siddiq and Santos, 2022; Pearce et al., 2022; He et al., 2024).
-Such prior benchmarks are composed of security-specific scenarios and are sourced from three resources, i.e., MITRE documentation (CWE-MIRTE, 2022), CodeQL repository (CodeQL, 2024), and scenarios manually crafted by the authors.
-We first combined all three existing benchmarks (Siddiq and Santos, 2022; Pearce et al., 2022; He et al., 2024), removed duplicates, and excluded the scenarios derived from MITRE examples, as Secure-Instruct uses MITRE documents as seeds for LLMs to generate fine-tuning datasets.
-Then, we followed a similar process as prior work He et al.
-(2024) to design 36 new security-specific scenarios covering 27 CWEs by adapting examples from the CodeQL repository.
-In the end, our crafted CWEBench consists of 93 security-specific scenarios covering 44 CWEs in six languages.
-Second, we include a recent benchmark CWEval (Peng et al., 2025), which uses test cases for automated verification of both functional correctness and security without relying on static security checkers.
-In addition, CWEval provides manual-crafted comprehensive specifications that describe complex user intentions and logic, which means it is a much closer step to real-world applications.
-In total, CWEval consists of 119 security scenarios covering 31 CWE types in five programming languages.
-De-duplication between fine-tuning datasets and evaluation benchmarks.
-To prevent data leakage between the fine-tuning dataset and the two benchmark datasets, we performed de-duplication between training and evaluation following  Zhou et al.
-(2025).
-Specifically, we used MinHash with Locality-Sensitive Hashing (LSH) to detect potential duplicates.
-Code snippets were grouped based on similarity, and pairs with a Jaccard similarity above 0.7 were flagged as potential duplicates.
-After that we removed these flagged snippets from the fine-tuning dataset.
-5 Research Question Results
-In this section, we present our three RQs. For each RQ, we detail the motivation, approach, and results.
-RQ1: To what extent can Secure-Instruct enhance the capabilities of LLMs to generate secure code?
-Motivation.
-We assess LLMs’ capabilities to generate secure code after being fine-tuned by Secure-Instruct.
-Approach.
+These models are widely used in prior works on secure code generation, such as (he2023large; he2024instruction; wei2024magicoder) for a fair comparison.
 For each of the four studied LLMs, we experimented with two settings of Secure-Instruct: (1) applying security-specialized instruction tuning on the dataset of pairs of vulnerable and secure code, i.e., from Secure-Instruct(vul, secure); and (2) applying standard instruction tuning on the secure code snippets from Secure-Instructsecure.
 Table 1 summarizes the statistics of the resulting fine-tuning datasets using Secure-Instruct.
-For CWEBench, we define a Secure Ratio to measure LLM’s capability to generate secure code.
-Secure Ratio of each security scenario is the percentage of secure code among all the generated code.
-The secure ratio is defined in Equation (3) where $S$ denotes the set of all generated code snippets and $\text{Sec}(s)$ evaluates to true if snippet $s$ is identified as non-vulnerable according to a static analyzer.
-Thus, the secure ratio quantifies the proportion of generated code that is free from vulnerabilities.
-To ensure robust results, we generated 100 code snippets for each security-specific scenario.
+Attribute Secure-Instruct Secure-Instruct
+(vul, secure) secure
+# of Code Snippets 856 15,600
+# of Average LOCs 52.0 50.0
+# of CWEs 25 44
+# of CWE-language pairs 39 78
+Table 1: The statistics of the fine-tuning datasets used in Secure-Instruct.
+Details on Instruction-Tuning LLMs.
+We use the parameter-efficient technique, LoRA (hu2021lora), to instruction-tune the 7B LLMs.
+LoRA reduces the GPU memory usage and maintains a promising performance.
+For the 1B model, we apply full-weight instruction-tuning.
+All experiments are conducted on five Nvidia RTX A6000 GPUs with 48GB of memory.
+For full-weight instruction-tuning, we set the learning rate to 2e-5, while for LoRA fine-tuning, the learning rate is 3e-4.
+Each model is tuned in 2 epochs.
+In LoRA, we configure the LoRA alpha to 16 to control the scaling of LoRA layers and set the dropout rate to 0.05.
+Evaluation Benchmarks.
+We evaluate Secure-Instruct using two secure-code generation benchmarks: (1) CWEBench crafted and extended by us, and (2) CWEval (peng2025cweval), which enables validation of both functional correctness and security by test cases.
+First, we craft CWEBench by extending prior benchmarks for secure code generation (siddiq2022securityeval; pearce2022asleep; he2024instruction).
+Such prior benchmarks are composed of security-specific scenarios and are sourced from three resources, i.e., MITRE documentation (CWE), CodeQL repository (codeql), and scenarios manually crafted by the authors.
+We first combined all three existing benchmarks (siddiq2022securityeval; pearce2022asleep; he2024instruction), removed duplicates, and excluded the scenarios derived from MITRE examples, as Secure-Instruct uses MITRE documents as seeds for LLMs to generate fine-tuning datasets.
+We define a secure ratio to measure LLM’s capability in CWEBench.
+Secure ratio is the percentage of secure code identified by CodeQL among all the generated code.
+To ensure robust results, we generated 100 samples for each security-specific scenario with a temperature of 0.8.
 A piece of LLM-generated code is deemed secure code if CodeQL detects no vulnerabilities in the code.
-$\text{Secure Ratio}=\frac{\left|\{s\in S\mid\text{Sec}(s)\}\right|}{|S|}$ (4)
+Then, we followed a similar process as prior work (he2024instruction) to design 36 new security-specific scenarios covering 27 CWEs by adapting examples from the CodeQL repository.
+In the end, our crafted CWEBench consists of 93 security-specific scenarios covering 44 CWEs in six languages.
+Second, we include a recent benchmark CWEval (peng2025cweval), which uses test cases for automated verification of both functional correctness and security without relying on static security checkers.
+In addition, CWEval provides manually crafted comprehensive specifications that describe complex user intentions and logic, which means it is a much closer step to real-world applications.
+In total, CWEval consists of 119 security scenarios covering 31 CWE types in five programming languages.
 CWEval comes with three metrics, namely Func@k, Sec@k and Func-Sec@k to evaluate the security and functionality of generated code.
@@
 Func-Sec@1 concerns both functional correctness and security.
-Given $n$ generated candidates, the unbiased Pass@k is computed by Equation (5.
-It defines the unbiased Pass@$k$ metric, denoted as $\text{Metric}_{\theta}@k$, where $\theta\in\{\text{Sec},\text{Func},\text{Func-Sec}\}$.
-Given $n$ generated candidates, the metric computes the probability that at least one of the top-$k$ candidates satisfies the evaluation criterion $\theta$.
-Thus, $\text{Metric}_{\theta}@k$ calculate the percentage of generated outputs that pass test cases under the chosen evaluation criterion within the top-$k$ samples.
-$\text{Metric}_{\theta}@k\;=\;1-\frac{\binom{n-c_{\theta}}{k}}{\binom{n}{k}}\qquad\theta\in\{\text{Sec},\text{Func},\text{Func-Sec}\}$ (5)
-Attribute Secure-Instruct(vul, secure) Secure-Instructsecure
-# of Code Snippets 856 15,600
-# of Average LOCs 52.0 50.0
-# of CWEs 25 44
-# of CWE-language pairs 39 78
-Table 1: The statistics of the fine-tuning datasets used in Secure-Instruct.
-Model Lang. Pre-trained (%) Secure-Instruct (%)
-(vul,secure) secure
-CodeLlama- 7B C 841/1,900 (44.3) 1,106/1,900 (58.2) 1,206/1,900 (63.5)
-Go 341/900 (37.9) 729/900 (81.0) 707/900 (78.6)
-Java 511/1,400 (36.5) 632/1,400 (45.1) 759/1,400 (54.2)
-JS 937/1,800 (52.1) 1,235/1,800 (68.6) 1,294/1,800 (71.9)
-Python 1,172/2,200 (53.3) 1,410/2,200 (64.1) 1,542/2,200 (70.1)
-Ruby 621/1,100 (56.5) 877/1,100 (79.7) 988/1,100 (89.8)
-Total 4,423/9,300 (47.6) 5,989/9,300 (64.4) 6,496/9,300 (69.8)
-CodeGen2- 7B C 875/1,900 (46.1) 963/1,900 (50.7) 956/1,900 (50.3)
-Go 488/900 (54.2) 390/900 (43.3) 617/900 (68.6)
-Java 446/1,400 (31.9) 489/1,400 (34.9) 513/1,400 (36.6)
-JS 956/1,800 (53.1) 805/1,800 (44.7) 1,147/1,800 (63.7)
-Python 1,190/2,200 (54.1) 1,269/2,200 (57.7) 1,388/2,200 (63.1)
-Ruby 584/1,100 (53.1) 638/1,100 (58.0) 683/1,100 (62.1)
-Total 4,539/9,300 (48.8) 4,554/9,300 (49.0) 5,304/9,300 (57.0)
-Mistral- 7B C 1,019/1,900 (53.6) 1,142/1,900 (60.1) 1,142/1,900 (60.1)
-Go 395/900 (43.9) 693/900 (77.0) 721/900 (80.1)
-Java 482/1,400 (34.4) 454/1,400 (32.4) 704/1,400 (50.3)
-JS 1,005/1,800 (55.8) 1,174/1,800 (65.2) 1,168/1,800 (64.9)
-Python 1,198/2,200 (54.5) 1,390/2,200 (63.2) 1,431/2,200 (65.0)
-Ruby 695/1,100 (63.2) 847/1,100 (77.0) 1,008/1,100 (91.6)
-Total 4,794/9,300 (51.5) 5,700/9,300 (61.3) 6,174/9,300 (66.4)
-StarCoder- 1B C 1,097/1,900 (57.7) 1,306/1,900 (68.7) 1,256/1,900 (66.1)
-Go 513/900 (57.0) 689/900 (76.6) 738/900 (82.0)
-Java 549/1,400 (39.2) 560/1,400 (40.0) 549/1,400 (39.2)
-JS 1,100/1,800 (61.1) 1,083/1,800 (60.2) 1,259/1,800 (69.9)
-Python 1,152/2,200 (52.4) 1,413/2,200 (64.2) 1,583/2,200 (72.0)
-Ruby 739/1,100 (67.2) 849/1,100 (77.2) 890/1,100 (80.9)
-Total 5,150/9,300 (55.4) 5,900/9,300 (63.4) 6,275/9,300 (67.5)
-Table 2: The secure ratio for pre-trained and fine-tuned models using Secure-Instruct on the full evaluation scenarios, broken down by language.
-Results.
-Table 2 presents the results on CWEBench using Secure-Instruct.
+De-duplication between fine-tuning datasets and evaluation benchmarks.
+To prevent data leakage between the fine-tuning dataset and the two benchmark datasets, we performed de-duplication between training and evaluation following  (zhou2025lessleak).
+Specifically, we used MinHash with Locality-Sensitive Hashing (LSH) to detect potential duplicates.
+Code snippets were grouped based on similarity, and pairs with a Jaccard similarity above 0.7 were flagged as potential duplicates.
+After that, we removed these flagged snippets from the synthesized dataset.
+4.2 Evaluation Results
+In this section, we present our evaluation results in both secure and functional benchmarks.
+Results of CWEBench.
+We evaluated the pre-trained models and instruction-tuned models by Secure-Instruct.
+We also conducted a comparative analysis between Secure-Instruct and SafeCoder (he2024instruction).
+In particular, we reused the two released models, i.e., CodeLlma-7B and Mistral-7B, which were already instruction-tuned by SafeCoder.
+Table 2 presents the results on CWEBench.
+The highest secure ratio is highlighted in bold.
 Across all models, Secure-Instruct(vul, secure) and Secure-Instructsecure demonstrate significant improvements in secure code generation.
-Compared with the pre-trained models, Secure-Instruct achieved the highest secure ratio across all models and languages.
-Specifically, CodeLlama‑7B achieved the highest secure ratio among the four models when fine‑tuned with Secure-Instructsecure, improving by 22.2% (from 47.6% to 69.8%), compared to a 16.8% increase under Secure-Instruct(vul, secure).
-CodeGen2-7B is improved from 48.8% to 57.0% by Secure-Instructsecure and to 49.0% by Secure-Instruct(vul, secure).
-Mistral-7B achieved an improvement of 14.9% in secure ratio by Secure-Instructsecure.
-StarCoder-1B achieves a notable improvement, i.e., secure ratio rising from 55.4% to 67.5%.
-Across the four LLMs, Secure-Instruct is shown to improve them to generate more secure code, and Secure-Instructsecure shows consistently better results than Secure-Instruct(vul, secure).
-Table 3 presents the results on CWEval.
-Secure-Instruct achieves the highest Func-Sec@1 across all the studied LLMs.
-Notably, CodeLlama-7B and Mistral-7B show significant improvements, with increases of 14.0% and 7.8%, respectively, compared to their pre-trained versions.
-Overall, the LLMs tuned by Secure-Instruct exhibit better secure code generation than their pretrained counterparts on both CWEBench and CWEval, across multiple security metrics.
-On CWEBench, Secure-Instructsecure surpasses Secure-Instruct(vul, secure) by an average of 5.7% in secure ratio.
-On CWEval, the two perform comparably in terms of Func@1, Sec@1, and Func-Sec@1.
+Compared with the pre-trained models, Secure-Instruct achieved the highest secure ratio across all models.
+For example, CodeLlama‑7B achieved the highest secure ratio among the four models when instruction‑tuned with Secure-Instructsecure, improving by 46.6% (the secure ratio from 47.6% to 69.8%), compared to an increase 35.3% under Secure-Instruct(vul, secure).
+The secure ratio of CodeGen2-7B is improved from 48.8% to 57.0% of secure ratio by Secure-Instructsecure and to 49.0% by Secure-Instruct(vul, secure).
+Mistral-7B achieved an increase of 28.9% by Secure-Instructsecure.
+StarCoder-1B also achieves a notable improvement, i.e., secure ratio rising from 55.4% to 67.5%.
+Across the four LLMs, Secure-Instruct improves them to generate more secure code, and Secure-Instructsecure shows consistently better results than Secure-Instruct(vul, secure).
+Compared to SafeCoder, our two models consistently outperform the SafeCoder models, with +15.2% in CodeLlama-7B and +9.9% in Mistral-7B using Secure-Instructsecure.
+At the CWE category level, CodeLlama‑7B instruction tuned with Secure-Instructsecure achieves the highest or ties for highest secure ratio in 29 of 44 CWEs, outperforming both its pre-trained baseline and the SafeCoder-tuned variant.
+Similarly, Mistral‑7B with Secure-Instructsecure leads in 25 of 44 CWEs relative to its pretrained and SafeCoder‑tuned counterpart.
+Due to the limited space, we put all detailed analyses in our replication package.
+These results demonstrate that Secure-Instructsecure significantly enhances secure code generation, outperforming the SOTA approach, SafeCoder.
+Code Llama- 7B Code Gen2- 7B Mistral- 7B Star Coder- 1B
+Pre-trained 47.6% 48.8% 51.5% 55.4%
+SafeCoder 60.6% ($\uparrow$ 27.3%) NA 60.4% ($\uparrow$ 17.3%) NA
+Secure-Instruct (vul,secure) 64.4% ($\uparrow$ 35.3%) 49.0% ($\uparrow$ 0.4%) 61.3% ($\uparrow$ 19.0%) 63.4% ($\uparrow$ 14.4%)
+Secure-Instruct secure 69.8% ($\uparrow$ 46.6%) 57.0% ($\uparrow$ 16.8%) 66.4% ($\uparrow$ 28.9%) 67.5% ($\uparrow$ 21.8%)
+Table 2: The secure ratio for pre-trained and instruction-tuned models using Secure-Instruct and SafeCoder on CWEBench.
+$\uparrow$ and $\downarrow$ indicate relatively change from pre-trained models.
+NA means the model not released by SafeCoder.
+Results of CWEval.
+Table 3 presents the results on the benchmark CWEval.
+Similar to CWEBench, Secure-Instruct achieves the highest Func-Sec@1 across all evaluated LLMs.
+Notably, CodeLlama-7B and Mistral-7B show significant improvements, with increases of 157.3% and 46.4%, compared to their pre-trained versions.
+For Sec@1, CodeLlama-7B, CodeGen2-7B, and StarCoder-1B achieved the highest secure ratio using Secure-Instruct, with 33.7%, 27.9%, and 16.4%, respectively.
+The results of Func@1 are put in our replication package due to limited space.
+Interestingly, SafeCoder exhibits a performance drop on CWEval.
+For example, SafeCoder’s CodeLlama-7B achieves only 7.1% Func-Sec@1, which is lower than the pre-trained model.
+A similar decline is observed in SafeCoder’s Mistral-7B.
+These findings are consistent with the observations reported in the study (peng2025cweval).
+Overall, LLMs with Secure-Instruct consistently outperform both their pre-trained versions and SafeCoder models in secure code generation across CWEBench and CWEval.
+CodeLlama-7B CodeGen2-7B Mistral-7B StarCoder-1B
+Sec@1 Func-Sec@1 Sec@1 Func-Sec@1 Sec@1 Func-Sec@1 Sec@1 Func-Sec@1
+Pre-trained 17.8% 8.9% 17.9% 7.4% 20.4% 12.5% 16.2% 6.1%
+SafeCoder 17.4% ($\downarrow$ 2.2%) 7.1% ($\downarrow$ 20.2%) NA NA 17.4% ($\downarrow$ 14.7%) 11.5% ($\downarrow$ 8.0%) NA NA
+Secure-Instruct (vul,secure) 31.2% ($\uparrow$ 75.3%) 22.9% ($\uparrow$ 157.3%) 17.7% ($\downarrow$ 1.1%) 7.2% ($\downarrow$ 2.7%) 25.5% ($\uparrow$ 25.0%) 17.9% ($\uparrow$ 43.2%) 16.4% ($\uparrow$ 1.2%) 7.4% ($\uparrow$ 21.3%)
+Secure-Instruct secure 33.7% ($\uparrow$ 89.3%) 22.9% ($\uparrow$ 157.3%) 16.4% ($\downarrow$ 8.4%) 7.5% ($\uparrow$ 1.4%) 27.9% ($\uparrow$ 36.8%) 18.3% ($\uparrow$ 46.4%) 15.5% ($\downarrow$ 4.3%) 7.4% ($\uparrow$ 21.3%)
+Table 3: Sec@1 and Func-Sec@1 results on CWEval for pre-trained LLMs and LLMs tuned by SafeCoder and Secure-Instruct (ours).
+Results of Functional Benchmarks.
+We evaluated our approaches with functional benchmarks, i.e., HumanEval and MBPP, two widely used benchmarks for assessing code generation in Python.
+For each task, we generate 40 samples and assess LLMs’ functionality using the Pass@k metric, including Pass@1.
+Pass@k is a standard metric for evaluating code generation, which represents the probability that at least one of the $k$ generated candidates for a given problem is correct.
+Table 4 presents the Pass@1 (Pass@10 is available in our replication package) on the HumanEval and MBPP benchmarks for pre-trained LLMs, instruction-tuned LLMs by SafeCoder and Secure-Instruct.
+For CodeLlama‑7B, using Secure-Instructsecure achieves the highest Pass@1 on HumanEval at 35.6%, outperforming both the pre-trained and SafeCoder versions.
+For Mistral‑7B, the model using the Secure-Instruct(vul, secure) achieves the highest score on MBPP at 30.7%.
+Codegen2-7B achieved the highest Pass@1 using Secure-Instructsecure, with 16.3% on HumanEval and 22.8% on MBPP.
+Overall, most models show performance improvements on both datasets, except for StarCoder-1B.
+CodeLlama-7B CodeGen2-7B Mistral-7B StarCoder-1B
+HumanEval MBPP HumanEval MBPP HumanEval MBPP HumanEval MBPP
+Pre-trained 25.5% 28.3% 15.3% 20.4% 23.0% 24.7% 11.7% 13.7%
+SafeCoder 32.0% ($\uparrow$ 25.5%) 33.7% ($\uparrow$ 19.1%) NA NA 31.3% ($\uparrow$ 36.1%) 29.3% ($\uparrow$ 18.6%) NA NA
+Secure-Instruct (vul,secure) 34.7% ($\uparrow$ 36.1%) 36.0% ($\uparrow$ 27.2%) 8.1% ($\downarrow$ 47.1%) 15.4% ($\downarrow$ 24.5%) 31.6% ($\uparrow$ 37.4%) 30.7% ($\uparrow$ 24.3%) 6.7% ($\downarrow$ 42.7%) 12.2% ($\downarrow$ 10.9%)
+Secure-Instruct secure 35.6% ($\uparrow$ 39.6%) 34.7% ($\uparrow$ 22.6%) 16.3% ($\uparrow$ 6.5%) 22.8% ($\uparrow$ 11.8%) 25.9% ($\uparrow$ 12.6%) 27.1% ($\uparrow$ 9.7%) 10.9% ($\downarrow$ 6.8%) 17.7% ($\uparrow$ 29.2%)
+Table 4: Pass@1 results on HumanEval and MBPP for pre-trained LLMs and LLMs tuned by SafeCoder-tuned and Secure-Instruct.
+5 Discussions
+The accuracy of CodeQL.
+In CWEBench, the ground-truth of security is based on CodeQL, which is a static tool and may produce false positives.
+To address this concern, we randomly sampled 291 code snippets labeled as secure by CodeQL from four models and manually verified their security.
+All sampled snippets were confirmed to be truly secure.
+To further support our conclusion, we used an additional benchmark, CWEval, which evaluates security through test cases.
+The similarity between Secure-Instruct’s instruction-tuning datasets and our evaluation benchmark.
+We performed a similarity analysis between the instruction-tuning data and benchmark data.
+Specifically, we first extracted all the secure code examples from the CodeQL repository for the 44 CWEs of our benchmark.
+We compare each LLM-generated secure code with the secure code examples from the CodeQL repository under the same CWE.
+We then calculated the cosine similarity of each pair using TF-IDF embeddings (sparck1972statistical).
+Figure 5 depicts the average cosine similarity score for the data generated by Secure-Instructsecure, which is 0.17, while the average score for data generated by Secure-Instruct(vul, secure) is 0.23.
+Secure-Instructsecure’s values align with findings from previous work (wei2024magicoder) on synthesizing instruction-tuning data using LLMs, which reported similarity scores typically ranging between 0.1 and 0.2 in the previous synthesis techniques.
+Importantly, this result indicates that the improvements from Secure-Instruct are not simply due to the inclusion of data from the same distribution of the CodeQL repository.
+Instead, the generated data provides additional diversity for the instruction-tuning dataset.
+Figure 4: Cosine similarities between Secure-Instruct’s instruction-tuning dataset and the secure code examples by CodeQL (i.e., our evaluation benchmark).
+Diversity of Secure-Instruct’s synthetic datasets.
+We followed a similar procedure to compute the pair-wise similarity for Secure-Instruct(vul, secure) and Secure-Instructsecure using TF-IDF.
+The similarity of Secure-Instruct(vul, secure) is 0.17.
+The similarity of Secure-Instructsecure is 0.16.
+Results show that datasets from both approaches are basically diverse.
+Low cost of Secure-Instruct in generating instruction-tuned datasets.
+Secure-Instruct generates high-quality instruction-tuning data with low cost and no human effort.
+Using ChatGPT (GPT-4o-2024-08-06), the total cost for generating all secure code snippets is USD 114.09.
+Similarly, the cost is USD 109.74 for Claude.
+The cost per CWE-language is USD 2.87.
+In contrast, SafeCoder (he2024instruction) faces scalability challenges, filtering 145M commits down to just 1,200 code pairs, and requiring manual selection of 465 high-quality examples.
+Secure-Instruct eliminates such overhead, offering a faster, cheaper, and automated alternative for generating secure and reliable datasets.
+Code Llama- 7B Code Gen2- 7B Mistral- 7B Star Coder- 1B
+Secure-Instruct(vul, secure) 64.4% 49.0% 61.3% 63.4%
+Secure-Instructsecure 61.9% 49.1% 62.0% 66.2%
+Table 5: The secure ratios of Secure-Instruct(vul, secure) and Secure-Instructsecure with the same size (865) of training dataset.
 Secure-Instruct(vul, secure) v.s.
 Secure-Instructsecure.
-As Secure-Instructsecure utilizes a much larger fine-tuning dataset (15.6K v.s.
-856), it is not surprising that Secure-Instructsecure outperforms Secure-Instruct(vul, secure).
-To understand more about the comparison, we down‑sampled the Secure-Instructsecure dataset to 865 examples, while ensuring samples covering across all CWE-language, to match the size of the Secure-Instruct(vul, secure) dataset.
+Secure-Instructsecure uses a much larger dataset than Secure-Instruct(vul, secure) (15.6K vs.
+856), so its better performance is expected.
+To compare fairly, we downsampled Secure-Instructsecure to 865 examples, covering all CWE-language pairs.
 Table 5 shows the comparison results across the four LLMs.
@@
 This shows that Secure-Instructsecure still generates highly effective synthetic data without being driven by vulnerable code first.
-It also indicates that there exist duplicates in the data by Secure-Instructsecure and opportunities for future improvement.
-Summary of RQ1: Secure-Instructsecure significantly improves the studied LLMs in secure code generation, achieving an average increase of over 10%.
-When using tuning datasets of the same size, Secure-Instructsecure and Secure-Instruct(vul, secure) achieve comparable performance.
-Model Metrics Pre-trained SafeCoder Secure-Instruct
-(vul,secure) secure
-CodeLlama- 7B Func@1 22.1% 14.9% 37.9% 34.9%
-Sec@1 17.8% 17.4% 31.2% 33.7%
-Func-Sec@1 8.9% 7.1% 22.9% 22.9%
-CodeGen2- 7B Func@1 20.7% NA 20.0% 20.9%
-Sec@1 17.9% NA 17.7% 16.4%
-Func-Sec@1 7.4% NA 7.2% 7.5%
-Mistral- 7B Func@1 30.1% 21.7% 34.4% 31.3%
-Sec@1 20.4% 17.4% 25.5% 27.9%
-Func-Sec@1 12.5% 11.5% 17.9% 18.3%
-StarCoder- 1B Func@1 15.3% NA 13.8% 12.7%
-Sec@1 16.2% NA 16.4% 15.5%
-Func-Sec@1 6.1% NA 7.4% 7.4%
-Table 3: Func@1, Sec@1 and Func-Sec@1 results on CWEval for the pre-trained LLMs, SafeCoder fine-tuned LLMs, and Secure-Instruct (ours) fine-tuned LLMs. NA means Model not released by SafeCoder.
-Models Secure-Instruct(vul, secure) Secure-Instructsecure
-CodeLlama-7B 5,989/9,300 (64.4) 5,758/9,300 (61.9)
-CodeGen2-7B 4,554/9,300 (49.0) 4,562/9,300 (49.1)
-Mistral-7B 5,700/9,300 (61.3) 5,770/9,300 (62.0)
-StarCoder-1B 5,900/9,300 (63.4) 6,157/9,300 (66.2)
-Table 4: The secure ratios of Secure-Instruct(vul, secure) and Secure-Instructsecure with the same size (865) of training dataset.
+6 Conclusion
+We presented Secure-Instruct, a novel framework for improving secure code generation by leveraging LLMs to synthesize instruction-tuning datasets.
+Secure-Instruct overcomes the key limitations of the SOTA technique, particularly addressing dataset imbalance, inefficient data curation, and the significant manual effort required.
+Unlike existing approaches, Secure-Instruct can be easily adapted to support new CWEs.
+We also introduce CWEBench, the largest benchmark for secure code generation.
+Our evaluation shows that Secure-Instruct significantly improves the performance of four LLMs, achieving an average improvement of 28.5% on CWEBench and a 140.8% improvement of Func-Sec@1 on CWEval.
+Results show that Secure-Instruct outperforms the SOTA method, demonstrating both scalability and efficiency.
+Appendix A Appendix A: Prompt Template in Data Synthesizing
+In our study, we leverage two schemes to synthesize data by vulnerable-code driven for Secure-Instruct(vul, secure) and secure-code driven for Secure-Instruct(vul, secure).
+In Secure-Instruct(vul, secure), Figure 5 shows the prompt template that we first prompt LLMs to generate vulnerable code snippets.
+Then another prompt is used for vulnerable code fixing, shown in Figure 6.
+In Secure-Instructsecure, the prompt shown in Figure 7 is similar to the first step of Secure-Instruct(vul, secure), and the only change is that we directly ask LLMs to generate secure code snippets.
+Prompt Template for Vulnerable Code Generation of Secure-Instruct(vul, secure) Task: Please generate an easily understandable vulnerable code snippet based on the following description of CWE.
+(Please consider the diversity of the vulnerable code examples for each generation.) CWE: {<CWE_OVERALL_DESCRIPTION>} A Vulnerable Code Example: This is {<LANGUAGE>} language example.
+{<CODE>}
+Explanation of the Example: {<EXPLANATION>} Can you generate a vulnerable code example for {<TARGET_LANGUAGE>} language?
+Figure 5: Prompt template used for vulnerable code generation in Secure-Instruct(vul, secure).
+Prompt Template for Fixing Vulnerable Code of Secure-Instruct(vul, secure) Task: Please fix the vulnerable code based on the following description of CWE: CWE: {<CWE_OVERALL_DESCRIPTION>} The Vulnerable Code:
+{<GENERATED_VULNERABLE_CODE>}
+Can you FIX the code for {<LANGUAGE>} ?
+Please make sure the code is secure to the CWE mentioned above and runnable.
+Task: Please generate an easily understandable secure code snippet based on the following description of CWE.
+(Please consider the diversity of the vulnerable code examples for each generation.)
+Figure 6: Prompt template used for vulnerable code fixing in Secure-Instruct(vul, secure).
+Prompt Template for Secure Code Generation of Secure-Instructsecure Task: Please generate an easily understandable secure code snippet based on the following description of CWE.
+(Please consider the diversity of the secure code examples for each generation.) CWE: {<CWE_OVERALL_DESCRIPTION>} A Vulnerable Code Example: This is {<LANGUAGE>} language example.
+{<CODE>}
+Explanation of the Example: {<EXPLANATION>} Can you generate a secure code example for {<TARGET_LANGUAGE>} language?
+Figure 7: Prompt template used for secure code generation in Secure-Instructsecure.
+Prompt Template for Instruction Generation Create a single, very short (maximum two sentences) non-detailed functionality description that could be used as a prompt to generate either of the code snippets below… {<The_Non-vulnerable_code_snippet>}
+Figure 8: Prompt template used for instruction generation.
+Appendix B Appendix B: Distribution of CWEBench
+Figure 9 shows the distribution of evaluated scenarios for languages and CWEs.
+Among 93 scenarios, 19 are for C, 9 are for Go, 14 are for Java, 18 are for JavaScript, 22 are for Python, and 11 are for Ruby.
+These scenarios cover 44 distinct CWEs, enabling a comprehensive evaluation.
+Figure 9: The distribution of languages and CWEs in CWEBench.
+Appendix C Appendix C: Detailed results of CWEBench
+Table 6 breaks down the detailed results on CWEBench of CodeLlama-7B and Mistral-7B in their pre-trained, SafeCoder, and Secure-Instruct, on each CWE in our evaluation benchmark.
+The highest secure ratio for each CWE is highlighted in bold.
+In the scenarios whose CWEs are covered in both approaches (the top section of the table), our analysis shows that CodeLlama-7B fine-tuned with Secure-Instructsecure achieved the highest overall secure ratio (70.6%), outperforming SafeCoder by 2.3% in secure ratio, Mistral-7B fine-tuned with Secure-Instructsecure achieved 68.9% secure ratio, with a 3.7% higher than SafeCoder.
+In the generalization evaluation (i.e., with the full security-specific scenarios of our benchmark), Secure-Instructsecuredemonstrates strong generalization capabilities across a broader range of CWEs for both CodeLlama-7B and Mistral-7B.
+Specifically, CodeLlama-7B with Secure-Instructsecure significantly outperforms SafeCoder, achieving a 69.2% secure ratio, while Mistral-7B achieves a 66.4% secure ratio.
+Table 6: The results of secure ratio in CodeLlama-7B and Mistral-7B.
+The top section lists CWEs that are included in the fine-tuning dataset of SafeCoder, while the bottom lists CWEs that are not covered in the SafeCoder dataset.
 CWE CodeLlama-7B Mistral-7B
-Pre-trained SafeCoder Secure-Instruct Pre-trained SafeCoder Secure-Instruct
-Secure (Vul, Secure) Secure (Vul, Secure)
+Pre-trained SafeCoder Secure-Instruct Secure Secure-Instruct (Vul, Secure) Pre-trained SafeCoder Secure-Instruct Secure Secure-Instruct (Vul, Secure)
 cwe-022 16.0 43.2 32.0 37.2 14.8 34.6 38.0 20.8
@@
 Total 47.6 60.6 69.8 64.4 51.5 60.4 66.4 61.3
-Table 5: The results of secure ratio in CodeLlama-7B and Mistral-7B.
-The top section lists CWEs that are included in the fine-tuning dataset of SafeCoder, while the bottom lists CWEs that are not covered in the SafeCoder dataset.
-RQ2: How is Secure-Instruct compared with the state-of-the-art (SOTA) approach in securing LLM-based code generation?
-Motivation.
-SafeCoder (He et al., 2024) represents the state-of-the-art in secure LLM-based code generation.
-It employs an automated pipeline to mine vulnerable code and corresponding fixes from a large corpus of open-source commits (145 million), followed by human intervention for post-processing and validation.
-The final dataset consists of (instruction, vulnerable code, secure code) triples used to fine-tune LLMs.
-In contrast, our proposed approach, Secure-Instruct , is a fully automated technique with zero human intervention.
-Given these fundamental differences, it is important to compare the performance of Secure-Instruct against SafeCoder to assess the effectiveness and scalability of our method.
-Approach.
-We further conducted a comparative analysis between Secure-Instructsecure and SafeCoder on CWEBench and CWEval.
-In particular, we used the two released models, i.e., CodeLlma-7B and Mistral-7B, which were already fine-tuned by SafeCoder.
-In CWEBench, there are 44 CWE types.
-However, the training dataset of SafeCoder does not include all of them.
-SafeCoder was fine-tuned on a dataset containing 23 CWEs, whereas Secure-Instruct was fine-tuned on a dataset covering 44 CWEs.
-There are 20 CWEs covered by both Secure-Instruct and SafeCoder.
-We used the same configuration as described in the SafeCoder paper, namely a temperature of 0.8 and 100 samples for each security-specific scenario.
-Results.
-Table 3 shows that Secure-Instruct achieves higher Func-Sec@1 on CWEval than SafeCoder for both CodeLlama‑7B and Mistral‑7B.
-Notably, SafeCoder exhibits a performance drop on CWEval.
-For example, SafeCoder’s CodeLlama-7B achieves only 7.1% Func-Sec@1, which is lower than the pre-trained model.
-A similar decline is observed in SafeCoder’s Mistral-7B.
-These findings are consistent with the observations reported in the study Peng et al.
-(2025).
-Table 5 breaks down the detailed results on CWEBench of CodeLlama-7B and Mistral-7B in their pre-trained, SafeCoder, and Secure-Instructsecure, on each CWE in our evaluation benchmark.
-The highest secure ratio for each CWE is highlighted in bold.
-In the scenarios whose CWEs are covered in both approaches (the top section of the table), our analysis shows that CodeLlama-7B fine-tuned with Secure-Instructsecure achieved the highest overall secure ratio (70.6%), outperforming SafeCoder by 2.3% in secure ratio, Mistral-7B fine-tuned with Secure-Instructsecure achieved 68.9% secure ratio, with a 3.7% higher than SafeCoder.
-In the generalization evaluation (i.e., with the full security-specific scenarios of our benchmark), Secure-Instructsecure demonstrates strong generalization capabilities across a broader range of CWEs for both CodeLlama-7B and Mistral-7B.
-Specifically, CodeLlama-7B with Secure-Instructsecure significantly outperforms SafeCoder, achieving a 9.2% improvement, while Mistral-7B achieves a 6.0% increase.
-At the CWE category level, CodeLlama‑7B fine‑tuned with Secure-Instructsecure achieves the highest or ties for highest secure ratio in 29 of 44 CWEs, outperforming both its pretrained baseline and the SafeCoder‑fine‑tuned variant.
-Similarly, Mistral‑7B with Secure-Instructsecure leads in 25 of 44 CWEs relative to its pretrained and SafeCoder‑fine‑tuned counterpart.
-These results demonstrate that Secure-Instructsecure significantly enhances secure code generation, outperforming the SOTA approach, SafeCoder.
-To assess statistical significance, we conducted Fisher’s exact tests comparing the number of secure code snippets generated by the pre-trained model and the model fine-tuned with Secure-Instructsecure for each CWE.
-CWEs with statistical significance (p<<0.05) are highlighted in bold, such as CWE-089.
-Notably, more than half of the CWEs in Secure-Instructsecure exhibit statistical significance, further demonstrating the robustness of our approach in enhancing secure code generation.
-Summary of RQ2: Safecoder encounters a degradation in CWEval benchmark.
-Secure-Instructsecure significantly outperforms SafeCoder with 9.2% and 6.0% improvements across CWEBench.
-Even for those CWEs that are part of SafeCoder’s fine-tuning dataset, Secure-Instructsecure achieves 2.3% and 3.7% higher secure ratios in CodeLlama-7B and Mistral-7B, respectively.
-RQ3: Can the LLM fine-tuned by Secure-Instruct preserve the functional correctness in code generation?
-Motivation.
-A critical concern in fine-tuning LLMs in secure code generation is understanding how fine-tuning impacts the functionality of the generated code.
-While fine-tuning with secure code examples aims to improve the security aspects of the output, it may compromise the model’s ability to produce functionally correct code.
-Approach.
-The benchmark, CWEval, evaluates the functionality of generated code by test cases with the metric Func@1.
-Besides, we evaluate the fine-tuned models on HumanEval and MBPP, two widely used benchmarks for assessing code generation in Python.
-For each task of HumanEval and MBPP, we generate 40 samples using both pre-trained and fine-tuned models and assess their functionality using the Pass@k metric, including Pass@1 and Pass@10.
-Pass@k is a standard metric for evaluating code generation, which represents the probability that at least one of the $k$ generated candidates for a given problem is correct.
-Results.
-Table 3 presents the Func@1 results for all models.
-All models, except for StarCoder-1B, have improvements after using Secure-Instruct.
-CodeLlama-7B has an over 10% increase in Func@1 compared with the pre-trained model.
-Table 6 presents the Pass@1 (Pass@10 is available in our replication package) on the HumanEval and MBPP benchmarks for pre-trained LLMs, fine-tuned LLMs by SafeCoder and Secure-Instruct.
-For CodeLlama‑7B, among the pretrained, SafeCoder‑fine‑tuned, Secure-Instruct(vul, secure)‑fine‑tuned, and Secure-Instructsecure‑fine‑tuned variants, Secure-Instructsecure-fine-tuned model achieves the highest Pass@1 on HumanEval at 35.6%.
-For Mistral‑7B, the Secure-Instruct(vul, secure)‑fine‑tuned model achieves the highest score on MBPP at 30.7%.
-Overall, most models show performance improvements on both datasets, except for StarCoder-1B.
-On CWEval, its Func@1 dropped from 15.3% to 13.8% with Secure-Instruct(vul, secure), and 12.7% with Secure-Instructsecure.
-On HumanEval, the Pass@1 score of StarCoder-1B dropped from 11.7% to 10.9% with Secure-Instructsecureand from 11.7% to 6.7% with Secure-Instruct(vul, secure).
-This decline may be due to the smaller size of StarCoder-1B compared to the other LLMs.
-Summary of RQ3: Secure-Instruct enhances secure code generation without sacrificing functional correctness.
-Across diverse LLMs and benchmarks, Secure-Instruct consistently outperforms both pre-trained LLMs and those fine-tuned by SafeCoder in generating functionally correct code.
-Model Benchmark Pre-trained SafeCoder Secure-Instruct
-(vul,secure) secure
-CodeLlama- 7B HumanEval 25.5% 32.0% 34.7% 35.6%
-MBPP 28.3% 33.7% 36.0% 34.7%
-CodeGen2- 7B HumanEval 15.3% NA 8.1% 16.3%
-MBPP 20.4% NA 15.4% 22.8%
-Mistral- 7B HumanEval 23.0% 31.3% 31.6% 25.9%
-MBPP 24.7% 29.3% 30.7% 27.1%
-StarCoder- 1B HumanEval 11.7% NA 6.7% 10.9%
-MBPP 13.7% NA 12.2% 17.7%
-Table 6: Pass@1 results on HumanEval and MBPP for the pre-trained LLMs, SafeCoder fine-tuned LLMs, and Secure-Instruct (ours) fine-tuned LLMs.
-6 Discussions
-The accuracy of CodeQL.
-In CWEBench, the ground-truth of security is based on CodeQL, which is a static tool and may produce false positives.
-To address this concern, we randomly sampled 291 code snippets labeled as secure by CodeQL from four models and manually verified their security.
-All sampled snippets were confirmed to be truly secure.
-Additionally, we used another benchmark, CWEval, which evaluates security through test cases, to further validate our results.
-The similarity between Secure-Instruct’s fine-tuning datasets and our evaluation benchmark.
-To clarify data leaks (i.e., overlap between fine-tuning dataset and benchmark), we performed a similarity analysis.
-Specifically, we first extracted all the secure code examples from the CodeQL repository for the 44 CWEs of our benchmark.
-Since most scenarios of the 44 CWEs are created based on the CodeQL repository, such secure code can be considered as ground-truth code for conducting such a similarity analysis.
-We compare each LLM-generated secure code with the secure code examples from the CodeQL repository under the same CWE.
-We then calculated the cosine similarity of each pair using TF-IDF embeddings (Sparck Jones, 1972).
-As there exist multiple secure examples for one CWE in the CodeQL repository, we selected the highest score, as the similarity score for an LLM-synthesized secure code snippet.
-Figure 6 depicts the average cosine similarity score for the data generated by Secure-Instructsecure is 0.17, while the average score for data generated by Secure-Instruct(vul, secure) is 0.23.
-Secure-Instructsecure’s values align with findings from previous work (Wei et al., 2024) on synthesizing fine-tuning data using LLMs, which reported similarity scores typically ranging between 0.1 and 0.2 in the previous synthesis techniques.
-In contrast, Secure-Instruct(vul, secure) exhibits a slightly elevated similarity score, exceeding this range.
-These results indicate that Secure-Instructsecure generates data with lower similarity compared to Secure-Instruct(vul, secure).
-Importantly, this result indicates that the improvements from Secure-Instruct are not simply due to the inclusion of data from the same distribution of the CodeQL repository.
-Instead, the generated data provides additional diversity for fine-tuning the dataset.
-Figure 4: Cosine similarities between Secure-Instruct’s fine-tuning dataset and the secure code examples by CodeQL (i.e., our evaluation benchmark).
-Diversity of Secure-Instruct’s synthetic datasets.
-We followed a similar procedure to compute the pair-wise similarity for Secure-Instruct(vul, secure) and Secure-Instructsecure using TF-IDF.
-The similarity of Secure-Instruct(vul, secure) is 0.17.
-The similarity of Secure-Instructsecure is 0.16.
-Results show that datasets from both approaches are basically diverse.
-Low cost of Secure-Instruct in generating fine-tuned datasets.
-We show that Secure-Instruct is able to generate high-quality fine-tuning data with low monetary cost and no human intervention.
-For example, when using ChatGPT (GPT-4o-2024-08-06), Secure-Instruct costs \$114.09 USD for generating all secure code snippets.
-Similarly, for Claude, the cost is \$109.74 USD for generating all secure code snippets.
-Compared to manually curated datasets, this synthesis method is highly cost-efficient.
-The cost per CWE-language is \$2.87 USD.
-Compared to our Secure-Instruct, prior work SafeCoder (He et al., 2024) imposes challenges on a high-efficient data curation process, i.e., a sharp reduction from hundreds of millions of code commits on open source repositories on GitHub to a total of 25,000 repositories that can be compiled and scanned by CodeQL for automatically detecting vulnerability-fixing commits, finally to merely 1,200 pairs of vulnerable and secure code examples.
-In addition, SafeCoder relies on manual effort to identify 465 high-quality code samples from the 1,211 samples produced by the automated pipeline.
-In contrast, Secure-Instruct demonstrates significant advancements in efficiency by reducing the time and cost required to construct high-quality datasets while maintaining a strong focus on generating reliable and secure code samples.
-CodeLlama- 7B CodeGen2- 7B Mistral- 7B StarCoder- 1B
-Secure-Instruct(vul, secure) 64.4% 49.0% 61.3% 63.4%
-Secure-Instructsecure 61.9% 49.1% 62.0% 66.2%
-Table 7: The secure ratios of Secure-Instruct(vul, secure) and Secure-Instructsecure with the same size (865) of training dataset.
-Secure-Instruct(vul, secure) v.s.
-Secure-Instructsecure.
-Secure-Instructsecure uses a much larger dataset than Secure-Instruct(vul, secure) (15.6K vs.
-856), so its better performance is expected.
-To compare fairly, we downsampled Secure-Instructsecure to 865 examples, covering all CWE-language pairs.
-Table 7 shows the comparison results across the four LLMs.
-Notably, the differences between the two are minimal, indicating essentially equivalent effectiveness.
-This shows that Secure-Instructsecure still generates highly effective synthetic data without being driven by vulnerable code first.
-7 Threats to Validity
-Internal Threats.
+Appendix D Appendix D: Limitations
 LLMs can produce different outputs for the same prompt.
-To prevent the huge variations, we set a large sample size to ensure the stabilization results.
-CWEBench relies on CodeQL to detect vulnerability.
+To prevent the huge variations, we set a large (100) sample size to ensure the stabilization of results.
+CWEBench relies on CodeQL to detect vulnerabilities.
 Such a static analyzer may produce false positives.
@@
 However, this process may fail to detect duplicates, leading to false negatives.
-External Threats.
 We follow prior work and choose LLMs with smaller parameter sizes to make fine-tuning feasible under limited computational resources.
-Fine-tuning larger models, such as ChatGPT-4o, would incur significantly higher costs.
+Instruction-tuning larger models, such as ChatGPT-4o, would incur significantly higher costs.
 As a result, we acknowledge that our findings may not fully generalize to larger LLMs, and we consider this a direction for future investigation.
@@
 To improve coverage, we expanded our security-sensitive prompts to incorporate additional CWEs beyond those included in earlier studies.
-Compared to prior work (Pearce et al., 2022; He and Vechev, 2023; He et al., 2024; Li et al., 2024a), our study evaluates a significantly greater number of CWEs, providing a more thorough assessment of generated code vulnerabilities.
+Compared to prior work (pearce2022asleep; he2023large; he2024instruction; li2024exploratory), our study evaluates a significantly greater number of CWEs, providing a more thorough assessment of generated code vulnerabilities.
 Future efforts should include broader CWEs.
@@
 Future work should expand the language set to strengthen the robustness of the results.
-8 Conclusion
-In this paper, we proposed Secure-Instruct, a novel framework to improve LLM for secure code generation by harnessing LLMs for generating instruction tuning datasets.
-Secure-Instruct overcomes the limitations of the state-of-the-art technique, particularly addressing dataset imbalance, inefficient data curation, and the significant manual effort required.
-Unlike existing approaches, Secure-Instruct can be easily adapted to support new CWEs.
-Additionally, we contribute CWEBench, the largest evaluation benchmark to date for secure code generation.
-Our evaluation shows that Secure-Instruct significantly improves the performance of four representative LLMs, achieving 14.3% more secure code on average on CWEBench and +5.3% Func-Sec@1 on average on CWEval.
-Compared to the state-of-the-art, Secure-Instruct demonstrates strong performance even with a smaller fine-tuning dataset, highlighting its scalability and efficiency.
-Finally, we show that LLMs fine-tuned with Secure-Instruct maintain comparable or improved functional correctness in generated code.
-Statements and Declarations
-Funding.
-This work was supported by the Fonds de recherche du Québec (Grant No.
-2024-NOVA-346499) [^1] https://repertoire.frq.gouv.qc.ca/offres/detailOffres.do;jsessionid=C333738136C878333C20227CC5EB4E1B?methode=afficher, the Natural Sciences and Engineering Research Council of Canada (NSERC) Discovery Grant (RGPIN-2019-07007), and the NSERC CREATE Grant (555406-2021).
-We gratefully acknowledge their support.
-Competing Interests. The authors have no financial or proprietary interests in any material discussed in this article.
-Data, Materials, and/or Code Availability. The data and code used in this study are available at https://zenodo.org/records/15019488.
-Authors’ Contributions.
-Junjie Li - Conceptualization, Methodology, Data collection, Analysis, Writing original draft.
-Fazle Rabbi - Material preparation, Data collection, Analysis.
-Bo Yang - Material preparation, Data collection, Analysis, Writing original draft.
-Jinqiu Yang - Conceptualization, Methodology, Writing review & editing, Supervision, Supervising experiment design and execution.
-Song Wang - Conceptualization, Methodology, Writing review & editing, Supervision, Supervising experiment design and execution.